### v1

# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

Abstract
Document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially in academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of academic papers sourced from arXiv, multiple strategies are proposed to generate automatically $1M$ QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal PDF understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
footnotetext: Equal contribution; Corresponding author
1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding [1, 2], where these models have demonstrated impressive capabilities in processing and generating human-like text. However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
To handle lengthy documents, current large models mainly fall their input into two separate modalities. One is the plain text modality. As shown in Fig. 1 (a), they focus only on the parsed pure text content and lack multimodal understanding of visual elements such as charts and figures within the documents. They usually rely on long-context LLMs [3, 4, 5], whose efficiency and accuracy decrease as the length of the document increases. Retrieval Augmented Generation (RAG) [6, 7, 8, 9] is also the main auxiliary means of these LLMs, but it requires the introduction of additional models and lacks the exploration of multimodal RAG.
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context/RAG LLMs for parsed pure text content. (b) Purely visual solution: VDU models for page-level encoding and feature interaction. (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
Another approach is to input documents in a purely visual modality, treating each page as an image. However, most current Visual Document Understanding (VDU) models models [10, 11, 12] only focus on visual QA on single-page documents. They usually rely on larger input resolution, resulting in more tokens, and thus have difficulty processing multi-page long documents. Recently, some multi-page VDU models can understand 8-page [13] or 20-page [14] documents. They usually encode each page separately and then perform page-level visual feature interactions such as concatenation [15, 16], as shown in Fig. 1 (b). However, more pages generate more visual tokens, which brings greater resource consumption to LLM. This makes them inefficient and unable to process longer documents.
Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong. Since most user queries are only related to a small part of the content in a long document, the sparse sampling can significantly remove redundant noise information. It encodes text paragraphs and diagrams in the parsed PDF document, utilizing both text and image representations to identify and extract the most relevant evidence in response to a user’s query. The sampled sparse evidence significantly reduces the number of input tokens of LLM and this process is independent of the length of the input documents. Moreover, the sparse sampler and LLM can be integrated in an end-to-end manner for training and inference, optimizing the performance of multimodal representation and question answering while improving time efficiency. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs. Another important characteristic is that it can naturally provide strong interpretability for the question answering.
In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for academic paper PDFs. The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model. The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies. We also provide a corresponding benchmark named PaperPDF.
The experimental results substantiate the effectiveness and efficiency of our approach on the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17]. Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
The main contributions of this paper are as follows:
  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient PDF question answering.
  • We propose a PDF multimodal question answering dataset (PaperPDF) with $1M$ QA pairs for training and $6k$ QA pairs for evaluation.
  • Our model significantly surpasses existing open-source models and proprietary products (by an average of $8.6\%$ on F1) on long multimodal PDF understanding.
Table 1: Comparison of various models for processing multi-page long documents.
Input modality Type Number of tokens Models
Plain text Long-context Linear increase LongLoRA [4], LongLLaMA [5], YaRN [3]
RAG w/o Linear increase Graph RAG [9], DISC-LawLLM [6], RAPTOR [19]
Purely visual Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], DocPedia [24], XC2-4KHD [25] and InternVL-V1.5 [11]
Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [26]
Text and images Unlimited-page w/o Linear increase PDF-WuKong (Ours)
2 Related Works
2.1 Document Understanding Datasets
Earlier document understanding datasets only focused on the NLP tasks such as summarization [27], and QA [28] of plain text. Meanwhile, there were several visual document datasets mainly aimed at text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34]. Recently, more datasets have been proposed for multimodal document QA across various scenarios. For example, DocVQA [35] and OCRVQA  [36] provide single-page QA data on books and business documents. ChartQA [37] and ChartX [38] focus on the visual reasoning for chart documents. ArXivQA [39] extracts scientific figure-caption data from the arxiv papers to enhance the academic ability of MLLMs. InfoVQA [40] contains many infographic documents which are a combination of textual, graphical and visual elements. However, these visual document datasets only define the single-page VQA task, and current MLLMs [41, 11] have achieved remarkable performance on it.
There are also some multi-page QA datasets [15, 17, 14] that require the model to understand the content relationship via multi-hop reasoning and capture the crucial information from multi-page documents. MP-DocVQA [14] extends DocVQA [35] by adding the context pages. DUDE [17] constructs multi-page QA data from multi-industry and multi-domain documents. Recently, DocGenome [42] was constructed as a scientific document benchmark. MM-NIAH [18] is a benchmark evaluating the capability of MLLMs to comprehend long multimodal documents, which requires the model to answer according to the key information scattered throughout the document. However, the answers in these datasets lack evidence and cannot provide reliable interpretability, especially for questions that require referring to multiple pieces of evidence in long documents.
2.2 Document Understanding Methods
Existing document understanding methods typically focus on either plain text or a limited number of document images. Methods that rely on pure text modality aim to process documents by first converting them into plain text through document parsing or Optical Character Recognition (OCR). They then employ efficient long-context mechanisms to handle long texts, such as sparse attention [4], memory networks [5], or position interpolation [3]. Besides, the methods based on retrieval-augmented generation (RAG) [6, 19, 9] also show impressive capabilities for long texts. While these approaches can integrate visual elements by including image captions or transforming images into natural language descriptions, they struggle with fine-grained understanding of visual information. This restricts their effectiveness in tasks requiring detailed interpretation of textual and visual components within documents.
Another solution is visual document understanding with purely visual input, treating each document page as an image. Many MLLMs such as UniDoc [20], mPLUG-DocOwl [21] and Vary [12] can perform this task in an OCR-free manner. Vary [12] employs an extra SAM-style [43] vision vocabulary specific to document and chart data, enabling direct encoding of entire pages with high compression ratios. Other researchers have advanced the understanding of high-resolution document pages by dividing input images into smaller patches, such as UReader [22], TextMonkey [10], and LLaVA-NeXT [23]. DocPedia [24] processes high-resolution images in the frequency domain via the DCT transformation. InternLM-XComposer2-4KHD [25] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration to capture more details. The reliance on high resolution results in a higher number of tokens and cannot be extended to multi-page documents.
Besides, there are several models specifically designed for multi-page documents. Hi-VT5 [14] and GRAM [16] are two professional models for multi-page QA. They combine image tokens of multiple pages through hierarchical transformer architecture [14] or global-local reasoning [16], being able to handle up to 20 pages. Fox [13] unifies all image tokens of up to 8 pages into a sequence to achieve multi-page QA. mPLUG-DocOwl2 [26] compress each high-resolution document image into 324 tokens, guided by low-resolution global visual features. These models encode each page separately and then perform page-level visual feature interactions. However, more pages generate more visual tokens, which brings greater resource consumption to LLM. This makes them inefficient and unable to process longer documents. Given this, we propose first to parse the lengthy document into interleaved text and image content and then perform sparse sampling in an end-to-end manner. Tab. 1 summarizes the attributes of each type of method.
3 Methodology
Figure 2: The overall structure of PDF-WuKong consists of a document parser, a sparse sampler and a large language model.
3.1 Overview
In order to achieve multimodal understanding of long PDF documents and alleviate the drawbacks of existing models that treat PDF documents as a single modality of plain text or images, we propose PDF-WuKong. The core motivation is that users’ queries are often only related to a small number of text blocks or diagrams in a long document. Therefore, to improve the accuracy and efficiency of MLLM, we design a sparse sampler integrated with a multimodal large language model in an end-to-end manner.
Specifically, our pipeline consists of three parts: a document parser, a sparse sampler and a large language model, as shown in Fig. 2. The document parsing stage first converts the input PDF document into machine-readable content of interleaved text and images, Then, the sparse sampler encodes the text blocks and images separately and caches their embeddings. When a user inputs a query, the most relevant content can be sampled using a simple similarity measure. Finally, the query and the sampled tokens are input into the LLM to generate the answer. Algorithm 1 shows the detailed steps of this method.
Algorithm 1 Inference pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$
2:Output: Generated answer $a$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Document Parsing
5:Parse the input document $D$ into text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Sparse Sampling
7:Encode all text blocks and images and cache all candidate vector embeddings:
$E_{T}=\{e_{T_{1}},e_{T_{2}},\dots,e_{T_{n}}\}\leftarrow\text{En\_T}(\{T_{1},T_{2},\dots,T_{n}\}),$
$E_{I}=\{e_{I_{1}},e_{I_{2}},\dots,e_{I_{m}}\}\leftarrow\text{En\_I}(\{I_{1},I_{2},\dots,I_{m}\})$
8:Encode the user query $q$:
$e_{q}\leftarrow\text{En\_T}(q)$
9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{E_{T},E_{I}\}$:
$S_{T}=\{\text{Sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
$S_{I}=\{\text{Sim}(e_{q},e_{I_{j}})\mid j=1,2,\dots,m\}$
10:Select the top-$k$ relevant text blocks and images:
$(T,I)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$
11:Stage 3: Answer Generation
12:Input the query $q$ and the selected tokens into the LLM:
$a\leftarrow\text{LLM}(q,(T,E_{I})_{top})$
13:Return the generated answer $a$.
Algorithm 2 Training pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$, ground truth answer $gt$
2:Output: Final loss function $\mathcal{L}_{\text{total}}$
3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
4:Stage 1: Data Preparing
5:Text blocks and images:
$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
6:Stage 2: Multimodal encoding
7:Encode the user query, positive and negative samples:
$e_{q}\leftarrow\text{En\_T}(q)$
$E_{T}=\{e_{T_{P}},e_{T_{N}}\}\leftarrow\text{En\_T}(\{T_{P},T_{N}\}),$
$E_{I}=\{e_{I_{P}},e_{I_{N}}\}\leftarrow\text{En\_I}(\{I_{P},I_{N}\}),$
8:Calculate the contrastive learning loss:
$\mathcal{L}_{\text{rep}}(e_{q},\{e_{T_{P}},e_{I_{P}}\},\{e_{T_{N}},e_{I_{N}}\})$
9:Stage 3: Output prediction of MLLM
10:Input the query, the positive text, and the positive image tokens from the shared image encoder En_I:
$a\leftarrow\text{LLM}(q,T_{P},e_{I_{P}})$
11:Calculate the cross-entropy loss:
$\mathcal{L}_{\text{QA}}(a,gt)$
12:Stage 4: Optimize model in an end-to-end manner
13:Update model parameters according to the joint loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}$
14:Return the final loss $\mathcal{L}_{\text{total}}$.
3.2 Document Parsing
Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure. By default, text blocks are organized into paragraphs, and all figures and tables are saved as images. These text and images are finally reorganized into an XML file in reading order. This process can be completed using existing open-source PDF parsing tools. During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
3.3 Sparse Sampling
For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems. The first is the problem of computing efficiency. The consumption of computing resources will increase dramatically. The second is the problem of inaccurate attention. Key information related to the user query is easily submerged by a large amount of irrelevant content. It is difficult for the model to accurately locate and extract important information in a huge token sequence. Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
For the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query. Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec. 4). We randomly select two text blocks and two images from the remaining text blocks and images as negative samples. Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$. An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query. The goal is to enable the model to capture the document content that is most relevant to the query. The contrastive learning loss is:
$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{{\text{sim}(e_{q},e_{i})}/\tau}}{e^{{\text{sim}(e_{q},e_{i})}/\tau}+\displaystyle\sum_{e_{j}\in E_{N}}e^{{\text{sim}(e_{q},e_{j})}/\tau}},$ (1)
where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples, respectively. $\tau$ is the temperature parameter that controls the scale of the similarity scores. $P$ is the number of positive samples. By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart. We align the feature space of the text encoder to the pre-trained vision encoder with frozen parameters. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
During the inference, we pre-encode all text blocks and images and cache all candidate vector embeddings. When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings. Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query. Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences. Moreover, the multimodal embedding cache further optimizes inference time.
3.4 Answer Generation
At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency. Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence. The inference process is shown in Algorithm 1.
Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process. Thus, the sparse sampler shares the same vision encoder with the MLLM. They can be integrated and trained in an end-to-end manner.
During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM. Besides, the query and instruction are also input into the LLM. Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth. Finally, the total optimization objective is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and question-answering. The training pipeline of PDF-WuKong is shown in Algorithm 2.
4 PaperPDF Construction
Figure 3: The construction process of PaperPDF based on single evidence and multiple evidence.
4.1 Overview
The reasons for constructing our dataset include the following two points. In most long document Q&A scenarios, the basis for an answer to a given question is typically composed of a single element or a combination of multiple elements, while the remaining information may act as noise, interfering with the reasoning of MLLMs. Additionally, existing document question-answering datasets are either limited to single-page formats or provide only page-level ground truth, which hinders the training of the sparse sampler in PDF-WuKong. Consequently, we propose a reliable method for generating high-quality question-answer pairs of long document and develop the dataset PaperPDF for training and evaluation based on this method.
Category Training Test
Single Text-only 249k 2939
Image-only 21K 212
Multi Image-text 250k 2566
Section 499k 255
Cross-paragraph 1.2k 118
Table 2: The statistic of PaperPDF. Text-only and Image-only indicate that the QA pairs are generated based on either a single text paragraph or an image extracted from the PDF. Image-text, Section, and Cross-paragraph denote that the QA pairs are generated from a paragraph and its corresponding references, an entire section, or non-contiguous paragraphs, respectively.
4.2 Detail of Construction
The construction process of PaperPDF are presented in Fig. 3. It can be divided into four steps: structured parsing, rule-based extraction, prompt construction, and filtering. We obtained 89k PDF documents from the arXiv repository as our document set $\mathbb{D}$. For each document $D$ in $\mathbb{D}$, we first employed Grobid to parse the document and extract its text chunks $\{T_{1},T_{2},\ldots,T_{m}\}$ and images chunks $\{I_{1},I_{2},\ldots,I_{n}\}$. Subsequently, predefined rules are used to randomly select chunks from the document, which may consist of text chunks $T_{i}$, chart chunks $I_{j}$ or a combination of both. The selected chunks will be input into commercial MLLM products according to different prompt templates and then, the question $Q$ related to the input chunks, along with the corresponding answer $A$, is subsequently generated. Notably, For the training set, we used Gemini for generation due to its free and rapid accessibility, while the high-performance GPT-4V was employed to construct the test set, ensuring the validity and robustness of the evaluation. Finally, we devise a set of rules to filter the generated training and testing sets automatically. The removing rules for the samples include too-short questions, too-long answers, non-English text, etc. Further manual checking is conducted on test set to ensure the reliability of the evaluation. PaperPDF consists of two types of QA pairs: Single-evidence and Multi-evidence.
Single-evidence QA pair means that the question can be answered based on a single text chunk or image chunk in the long document. It can be systematically categorized into two types: text-only and image-only. In text-only generation, a text chunk is input into MLLM, whereas in image-only generation, an image chunk is utilized. The generation process for Single-evidence QA pairs is relatively straightforward and cost-effective, primarily aimed at training the foundational multimodal understanding and question-answering capabilities of MLLMs.
Multi-evidence QA pairs. Unlike single-evidence QA pairs, Multi-evidence QA data exhibits greater complexity, both in terms of characteristics and generation process. The answers to such questions typically rely on multiple text chunks, image chunks, or any combination thereof in document. Multi-evidence QA pairs consist of Image-text QA pairs, Section QA pairs, and Cross-paragraph QA pairs. The Image-text QA pairs are generated based on a paragraph and its corresponding image references. These pairs require consideration of the interrelationship between the text and the image to ensure the answers capture multimodal information. Section QA pairs are designed to train the MLLM in integrating and understanding information within a section and are generated from all chunks within a section. Cross-paragraph QA pairs involve the most complex generation process. First, the entire document is input into the MLLM for paragraph-level semantic summarization, followed by the selection of semantically related text chunks. Finally, multiple related chunks are randomly selected and re-input into the MLLM to generate the final QA pairs. This type of data primarily focuses on training the sampler’s performance, enhancing its sparse sampling capability. Although the generation of Multi-evidence QA pairs is relatively complex and may require multiple MLLM calls, their presence significantly improves the performance of the sampler and strengthens the MLLM’s ability to comprehend cross-chunk information.
Totally, we generated 1.5M samples for training and 8K samples for testing. After filtering, the final dataset consists of 1M training data points and 6K test data points for subsequent training and evaluation. The statistics of PaperPDF are presented in Tab. 2.
Figure 4: Examples from PDF-WuKong and other proprietary products
5 Experiments
[…]
Model #param ANLS F1 ROUGE Token Latency GPT Acc
[…]
Closed Qwen [46] - 36.0 40.3 35.5 - - 78.1
Kimi [47] - 28.5 33.6 31.1 - - 74.7
Gemini pro [48] - 26.6 29.0 29.8 - - 67.9
ChatGLM [49] - 31.2 35.4 32.0 - - 73.5
[…]
5.1 Implementation Details
[…]
Table 4: Description of the input format used in the experiment
Symbol Description
* Input the parsed content of the pdf file.
[…]
$\dagger$ Input the entire page image.
$\ddagger$ Input the entire page image and OCR content.
5.2 Long PDF Understanding
To assess the effectiveness of our model in understanding long PDF documents, we conducted comprehensive experiments comparing it with both open-source models and commercial products on PaperPDF dataset.
Due to the limited availability of open-source models capable of handling this task, we firsly evaluate two such models that can process multi-page PDF documents. Considering that IXC2-VL [45] cannot understand such long image-text interleaved sequences, we also report the results with the input of pure OCR content. To further demonstrate the advantages of our multi-modal sampler, we demonstrate the evaluation of IXC2-VL with RAG where we feed the top 5 text paragraphs retrived by BGE-M3 [50]. The results on the PaperPDF dataset are reported in Tab. 3. The findings indicate that our model significantly outperforms existing open-source models across various evaluation metrics. Additionally, thanks to the introduction of the sparse sampler, the number of tokens that our model’s MLLM needs to process is substantially fewer than that of other models.
Moreover, we compared our model with several commercial products that allow users to input PDF documents and questions via web interfaces to generate answers. Considering the cost of manual testing, we randomly selected 50 PDFs from the PaperPDF dataset for evaluation, and the results are presented in Tab. 3. Our model clearly outperforms mainstream PDF question-answering products. Fig. 4 shows some examples from PDF-WuKong and other proprietary products. Our model can accurately retrieve the evidence needed to answer user questions and generate accurate answers based on relevant text segments or diagrams. Since these commercial solutions do not provide detailed code or technical reports, we could not obtain information about their parameter sizes or token counts, and uploading documents directly to their websites does not allow us to accurately assess the models’ latency. Finally, we evaluated the GPT accuracy of all open-source and proprietary models on this subset benchmark. The results show that our model has significant advantages in both textual similarity and semantic understanding.
To compare with more open-source large document models, considering that most of these models can only handle single-page documents, we constructed a subset of the PaperPDF benchmark containing only test samples with single evidence. Therefore, we provided all models with only the page containing the evidence as input; note that the pages are input in the form of images. These models and our PDF-WuKong have not been trained on such data. In this setting, our sparse sampler sampled all the content of that page. As shown in Tab.5, our model’s zero-shot capability on this subset is significantly better than other document large models.
Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on Single-Evidence Subset of PaperPDF. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model # param ANLS F1 ROUGE
[…]
5.3 Document-oriented VQA
To validate the generalization capability of our model to other document understanding scenarios, we conducted experiments on several benchmark datasets and compared PDF-WuKong with other representative models.
First, we evaluated the performance of PDF-WuKong and other open-source models on DocVQA [35], ChartQA [37], and InfoVQA [40], which are all single-page document datasets. As shown in Tab. 6, our model achieved leading performance compared to other open-source models. This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question answering tasks.
In addition, we assessed the performance of traditional specialized models and large-scale models on two existing multi-page document QA datasets. The experimental results, presented in Tab. 7, indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document large model, DocOwl2 [26]. Notably, on complex multi-page document datasets like DUDE [58], PDF-WuKong outperforms GPT-4V [59]. This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
Furthermore, we conducted zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18]. As shown in Tab. 8, our model uses the fewest parameters yet achieves the second-best performance. Although InternVL-V1-5-RAG [18] surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model. Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, unlike other models that experience significant declines. At a context length of 64K, PDF-WuKong achieves the best results, demonstrating its robustness in handling long-context multimodal inputs.
Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model Doc. Chart. Info.
Closed Gemini Pro [60] 88.1 74.1 75.2
GPT-4V [61] 88.4 78.5 75.1
[…]
Table 7: Performance comparison with other DocVLMs for multi-page document understanding. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
Model MP-DocVQA DUDE
[…]
Table 8: Performance comparison with other DocVLMs on MM-NIAH. The input approach aligns with the benchmark.
Model #param Overall 1K 4K 16K 64K
Emu2-Chat [55] 37B 8.8 38.9 18.2 0.0 0.0
VILA1.0-13b [69] 13B 15.7 41.9 33.2 8.6 0.1
llava-v1.6-13b [70] 13B 16.9 43.7 34.9 13.6 0.0
llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
InternVL1.5-RAG [18] 45B 46.1 59.5 50.1 44.9 39.3
PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
5.4 Ablation Study
To comprehensively evaluate the effectiveness of our proposed model components, we conducted ablation studies focusing on datasets, the impact of the sparse sampler, sampling strategies, and document length. Below, we present the findings from each of these experiments.
Datasets
We conducted experiments to evaluate the impact of training data size and composition on model performance. As shown in Tab. 9, increasing the amount of training data led to consistent improvements in the model’s accuracy on the PaperPDF benchmark. Specifically, the model trained with larger datasets demonstrated better comprehension and answer accuracy. Moreover, under the same data volume, training sets that included multi-evidence annotations outperformed those containing only single-evidence. This suggests that exposure to diverse and complex evidence during training enhances the model’s ability to handle queries requiring multiple pieces of information, thereby improving overall performance.
Table 9: Ablation study on dataset setting
Dataset ANLS F1 ROUGE
100 k 38.7 40.1 37.5
500 k 41.6 43.5 40.8
1 M 42.6 43.6 40.2
Impact of sparse sampler
To assess the effectiveness of the sparse sampler, we compared models trained with and without it. Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the overwhelming amount of irrelevant information. Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab. 10, by efficiently selecting the most relevant content for each query. Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately. This indicates that our integrated design allows the two components to mutually enhance each other, optimizing both the retrieval of pertinent information and the generation of accurate answers.
Table 10: Ablation study on the impact of sparse sampler
Sparse Sampler End-to-End ANLS F1 ROUGE
11.1 5.1 5.0
40.3 42.3 39.8
42.6 43.6 40.2
Sampling strategy
We explored different sampling strategies by varying the number of top N text paragraphs or figures selected by the sparse sampler. As shown in Tab. 11, setting N too low resulted in missing critical information necessary for accurately answering queries, thereby reducing performance. Conversely, a larger N introduced redundant information and increased computational and time costs without significant improvements in accuracy. This suggests that there is an optimal range for N that balances the inclusion of essential information with efficiency. Selecting the appropriate number of samples is crucial for maximizing performance while minimizing resource consumption.
Table 11: Ablation study of different sampling strategy
Sampling chunks ANLS F1 ROUGE Tokens
Top 1 39.20 38.28 35.26 1186
Top 3 42.09 43.06 39.69 1452
Top 5 42.59 43.63 40.19 1789
Top 10 43.01 44.22 40.67 2386
Top 15 43.19 44.57 42.08 2704
Top 20 43.42 45.02 42.30 3364
Figure 5: Ablation study of different document length
Document length
To understand the impact of document length on model performance and efficiency, we divided the test set into subsets based on the number of pages per document. Results in Fig. 5 demonstrate that our model’s performance and time efficiency remained relatively stable across documents of varying lengths. This stability indicates that the sparse sampler effectively reduces the input size to a manageable level, regardless of the original document length. In contrast, MLLMs without the sparse sampler were unable to handle long documents effectively; their performance and time efficiency deteriorated significantly as the document length increased. These findings highlight the robustness of our approach in processing long documents without sacrificing accuracy or incurring additional computational costs.
6 Conclusion
We have presented PDF-WuKong, a novel Multimodal Large Language Model that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images. By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length. We also constructed PaperPDF, a comprehensive dataset with 1 million question-answer pairs for training and 6,000 pairs for evaluation, specifically tailored for academic PDFs. Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on long multimodal PDF understanding tasks. Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
Appendix A Dataset Examples
Text-only Data Example in the Training Set Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model? Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig. 5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig. 5b). Answer 1: It can lead to an unfair advantage for the model. Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset. If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
Text-image Data Example in the Training Set Query: What is the best performing method for both detection and classification, according to the provided figure? Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1). Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches. Note: the -means the results were not reported by that method. Answer 1: Ours. Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks. This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
Section Data Example in the Training Set Query: What kind of information does the factor R i, j hold? Text:
RECOMMENDATION MODEL
In this section, we present our Joint Topic-Semantic-aware Matrix Factorization (JTS-MF) model for online social votings, in which social relationship, group a liation, and topic-semantic similarities are combined and taken into account for voting recommendation in a comprehensive manner.Motivated by Locally Linear Embedding [23] which tries to preserve the local linear dependency among inputs in the low-dimensional embedding space, we expect to keep inter-user and inter-voting topic-semantic similarities in latent feature space as well.To this end, in JTS-MF model, while the rating R i, j is factorized as user latent feature Q i and voting latent feature P j , we deliberately enforce Q i and P j to be dependent on their social-topic-semantic similar counterparts, respectively.e graphic model of JTS-MF model is as shown in Figure 5. Figure: Figure 5[Uncaptioned image] Caption: Fig. 5: Graphic Model of JTS-MF. Answer 1: The factor R i, j represents the rating that user i gave to voting j. Answer 2: The node labeled R i, j is a factor that represents the rating given by user i to a voting j. It is the core of the model, representing the essence of user interaction with the voting system. This node’s existence suggests that the model is primarily concerned with understanding and predicting user ratings for voting content.
Text-only Data Example in the Test Set Query: How does the distribution of output image transform tasks among N workers impact the efficiency of the final synchronization task? Text: The number of output image transform tasks equals the number of output images times the batch size. The tasks are executed by all N workers, such that each worker picks up an arbitrary task and executes it. The last output image transform task to finish also executes the final synchronization task, which frees the memory required for the output image transforms. Answer 1: The distribution allows for parallel processing, potentially speeding up the final synchronization task. Answer 2: Ensures faster task completion, synchronization efficiency relies on all tasks’ completion.
Text-image Data Example in the Test Set Query: What distinguishing features can be observed between ‘hit’ and ‘miss’ samples for real fingerprints in Fig. 5 across different benchmarks? Text: Images found in these benchmarks can be observed in Fig. 5 of Section V.As we can see, variability exists not only across modalities, but also within modalities.Moreover, it is rather unclear what features might discriminate real from spoofed images, which suggests that the use of a methodology able to use data to its maximum advantage might be a promising idea to tackle such set of problems in a principled way. Figure: Figure 5[Uncaptioned image] Caption: Fig. 5. Examples of hit and missed testing samples lying closest to the real-fake decision boundary of each benchmark. A magnified visual inspection on these images may suggest some properties of the problem to which the learned representations are sensitive. Answer 1: Hit’ samples for real fingerprints show clearer ridge patterns and quality compared to ‘miss’ samples throughout the benchmarks. Answer 2: Hit real fingerprints are clearer, more distinct; Miss real samples are blurred, less defined.
Only-image Data Example in the Test Set Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period? Figure: Figure 4[Uncaptioned image] Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour. A benchmark is counted for multiple techniques if two techniques are tied for the lead. Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4. Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
Section Data Example in the Test Set Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries? Text:
Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers.We create word clouds for the two best methods in Section 5.3.In this paper, we used an online tool called WordItOut5 to generate the word cloud representations.In all word clouds presented in this paper, a filter is used to display only the words with minimum frequency of 2.
Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method.This gives a sense of the content in those summaries.For contrast, we also generate a word cloud for the original news articles without the content of the generated summaries.Specifically, common words are first removed completely and then the word clouds are built with frequencies of surviving words.In essence, this shows what information remains apart from the generated summaries. The images clearly show a contrast of content.The summary wordcloud shows “earthquake” as its most prominent word.The image of the articles show less focus.If viewed alone, the reader would not quickly infer the gist of the original content.Similarly, Figure 4 represents “Lead” method.And here we see an even more stark difference.6. Comparing Twitter and News Media information. Figure: Figure 3[Uncaptioned image] Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right). Answer 1: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually. Answer 2: Visual contrast in word frequency highlights content focus differences.
Cross-paragraph Data Example in the Test Set Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms? Text:
Paragraph 7:
Figure 1: On the right: the current path for selecting some algorithm(s) based on their validity and performance for a given dataset.On the left: the proposed flip to better understand the space in which ‘good’ datasets exist for an algorithm.
Paragraph 17:
Section 2 describes the structure of the proposed method including its parameters and operators.
Paragraph 20:
In this section, the details of an algorithm that generates data for which a given function, or (equivalently) algorithm, is well-suited is described.This algorithm is to be referred to as “Evolutionary Dataset Optimisation” (EDO).
Figure: Figure 1[Uncaptioned image] Caption: Sample number of rows and columns 2. Sample columns and fill/trim values as needed. Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection. Answer 2: By employing the EDO approach, we can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use. This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy. As a result, the performance metrics of the algorithms improve significantly, leading to more valid outcomes and reliable interpretations. The focus on tailored dataset selection directly contributes to optimal algorithm efficiency, thereby validating the effectiveness of the EDO approach in practical applications.
Appendix B Prompt Engineering
Text-only QA Generation Prompt for the Training Set Task Definition: Generate 3 high quality academic questions and corresponding answers based on the provided text/image/table. Requirements: The questions should require very complex reasoning and global understanding and the answer should be detailed and can be answered by the material, do not add extra information. Remember it is only a paragraph rather than a whole passage so do not ask global question about the study. Generate question-answer pairs instead of multiple-choice questions. Do not generate global question about the paper such as main idea or abstract. You should use English. Expected Output: List your Q and A as: [Q1]: [A1]: [Q2]: [A2]: [Q3]: [A3]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea.
Figure 6: Text-only QA generation prompt for the training set.
Text-image QA Generation Prompt for the Training Set Task Definition: Generate 2 high-quality academic questions and corresponding answers based solely on a provided figure of a research paper, without relying on accompanying text for the questions. However, you may use the provided text to understand the figure. Ensure these questions demand complex reasoning, focusing exclusively on details within the figure that are critical for understanding. The answers must be detailed, directly reflecting the content in the figure and capable of being derived solely from the figure without extraneous information. Requirements: Thoroughly analyze the provided figure of the research paper, attaining a deep understanding of its contents, including but not limited to, the study’s objectives, methodologies, results, and conclusions. Develop 2 academic questions that: Must not mention the name of the figure directly or words like “figure”,“table” in the question. Demand complex reasoning and a comprehensive understanding of the whole figure. Must ask questions that can only be answered by the content in the figure, without reliance on textual information. Integrate knowledge from the entire figure, reflecting a nuanced understanding. Frame questions to elicit detailed and informed responses directly supported by the figure. Craft detailed answers for each question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: List your Q and A as: [Q1]: [Formulate the first academic question here, ensuring it requires complex reasoning and encompasses the entirety of the provided figure for a comprehensive understanding. Based solely on the figure, without referencing accompanying text. Must not mention the name of the figure directly. Must not mention words like “figure”, “table” exc.] [A1]: [Provide a detailed answer here, derived exclusively from the information within the provided figure, ensuring the response is thorough and precise without including extraneous details. Mention from which figure/table you get the answer.] [Q2]: [A2]: Here is the supplementary paragraph text for the figure:
Figure 7: Text-image QA generation prompt for the training set.
Section QA Generation Prompt for the Training Set Task Definition: Generate 3 high-quality academic questions and corresponding answers based on a section of a research paper. Requirements: Each question should generate two types of answers. The first answer should be concise, directly addressing the question with minimal wording. The second answer should include a “chain of thought” that provides a reasoning process and be detailed. You should use English. Generate questions that: Can be answered with simple reasoning and only require a global understanding. The question should be able to answered directly with the material. Do not include 2 or more subquestions in each question. Ensure that the concise answers can be provided using sentences or phrases that do not exceed 20 words in length. Remember it is only a paragraph rather than a whole passage so do not ask global question about the study. Generate question-answer pairs instead of multiple-choice questions. Do not generate global question about the paper such as main idea or abstract. Must not mention the name of the figure directly or words like “figure”, “table” exc in the question. Must ask questions that can only be answered by the content in the figure, without reliance on textual information. Craft the 2 answers for each question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: List your Q and A as: [Q1]: [Insert the academic question here, can be answered directly from the material] [A11]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words] [A12]: [Insert the detailed answer here, based solely on the given information. This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption. Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] [Q2]: [A21]: [A22]: [Q3]: [A31]: [A32]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea. Here is the supplementary paragraph text for the figure:
Figure 8: Section QA generation prompt for the training set.
Cross-paragraph Question Generation Prompt for the Training Set Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question. This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper. Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided). Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper. Note that your question should not directly include the “idx”s of the paragraphs. Expected Output: [Q]: [Your generated question based on the shared themes or information] Here are the selected paragraphs in the paper:
Cross-paragraph Answer Generation Prompt for the Training set Task Definition: Given some selected paragraphs from a research paper, each chosen based on their relevance (excluding the first 3 paragraphs), and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs. The question is crafted to encompass the themes or findings presented in the paragraphs of the chosen paragraphs, aiming for a comprehensive understanding and connection between these elements. Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material. Are comprehensive and cover all relevant aspects, as presented in the provided figure. Accurately reflect the information and insights offered by the research paper figure. Expected Output: [A1]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words.] [A2]: [Insert the detailed answer here, based solely on the given information. This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption. Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] The question is:
Figure 9: Cross-paragraph QA generation prompt for the training set.
Text-only Question Generation Prompt for the Test set Task Definition: Create 2 academic questions from a given research paper paragraph. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: No more than 30 words. Incorporate knowledge from the paragraph. Should be answered by text instead of one of the multiple choices. Elicit detailed responses supported by the text. Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
Text-only Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-only Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 10: Text-only type QA generation prompt for the test set.
Text-image Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on the provided figures and tables from a research paper. Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions. Focus on the paragraph,not the entire paper. If the paragraph lacks valid information,return ‘quit’. You should use English. Develop 2 questions that: No more than 30 words. Are specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables. Must not mention the label of the figure/table directly or use words like ‘from the figure/table’. Expected Output: [Q1]: [Q2]:
Text-image Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Always use English. Do not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Text-image Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 11: Text-image QA generation prompt for the test set.
Section Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on a section from a research paper. Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions. You should Always use English. Develop 2 questions that: No more than 30 words. Require an integration of information from all paragraphs and figures/tables in the section. Must not mention the label of the figure/table directly or use words like ‘from the figure/table’. Not based on common knowledge or assumptions not supported by the figures and tables. Expected Output: [Q1]: [Q2]:
Section Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: No more than one sentence, less than 20 words. Always use English. Do not infer or speculate on details not explicitly explained by the figures/tables. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence. [thinking procedure]: [A2]: answer2 here, no more than one sentence.
Section Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given. Requirements: The answers should be: Within a few keywords, less than 20 words. Comprehensive and cover all relevant aspects. Accurately reflect the paragraph’s information and insights. You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords. [thinking procedure]: [A2]: answer2 here, within a few keywords.
Figure 12: Section QA generation prompt for the test set.

### v2

@@
 Abstract
-Document understanding is a challenging task to process and comprehend large amounts of textual and visual information.
+Multimodal document understanding is a challenging task to process and comprehend large amounts of textual and visual information.
 Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task.
-However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially in academic papers.
+However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers.
 In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents.
@@
 The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model.
-To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of academic papers sourced from arXiv, multiple strategies are proposed to generate automatically $1M$ QA pairs along with their corresponding evidence sources.
-Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal PDF understanding, surpassing proprietary products by an average of 8.6% on F1.
+To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers.
+Multiple strategies are proposed to automatically generate 1.1 million QA pairs along with their corresponding evidence sources.
+Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1.
 Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
@@
 However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
-To handle lengthy documents, current large models mainly fall their input into two separate modalities.
-One is the plain text modality.
+To handle lengthy documents, current research in multimodal document understanding with LLMs primarily follows two mainstream technical routes.
+The first route is based on pure text modality understanding.
 As shown in Fig.
-1 (a), they focus only on the parsed pure text content and lack multimodal understanding of visual elements such as charts and figures within the documents.
-They usually rely on long-context LLMs [3, 4, 5], whose efficiency and accuracy decrease as the length of the document increases.
-Retrieval Augmented Generation (RAG) [6, 7, 8, 9] is also the main auxiliary means of these LLMs, but it requires the introduction of additional models and lacks the exploration of multimodal RAG.
+1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables).
+They then employ long-context LLMs [3, 4, 5] or utilize Retrieval Augmented Generation (RAG) techniques [6, 7, 8, 9] to process the textual content.
+The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
 Figure 1: Method comparison for long multi-page PDF document understanding.
 (a) Plain text solution: long-context/RAG LLMs for parsed pure text content.
-(b) Purely visual solution: VDU models for page-level encoding and feature interaction.
+(b) Pure vision solution: VDU models for page-level encoding and feature interaction.
 (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
-Another approach is to input documents in a purely visual modality, treating each page as an image.
-However, most current Visual Document Understanding (VDU) models models [10, 11, 12] only focus on visual QA on single-page documents.
-They usually rely on larger input resolution, resulting in more tokens, and thus have difficulty processing multi-page long documents.
-Recently, some multi-page VDU models can understand 8-page [13] or 20-page [14] documents.
-They usually encode each page separately and then perform page-level visual feature interactions such as concatenation [15, 16], as shown in Fig.
-1 (b).
-However, more pages generate more visual tokens, which brings greater resource consumption to LLM.
-This makes them inefficient and unable to process longer documents.
+The second route is based on pure visual modality understanding.
+As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks[10, 11, 12].
+However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues.
+This makes it challenging to efficiently process multi-page long documents.
+While some recent multi-page VDU models handle up to 8 pages [13] or 20 pages [14], they typically encode each page separately and perform page-level visual feature interactions such as concatenation [15, 16].
+Nevertheless, as the number of pages grows, computational resource consumption still escalates substantially, rendering these models inefficient for processing longer documents.
 Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong.
@@
 Another important characteristic is that it can naturally provide strong interpretability for the question answering.
-In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for academic paper PDFs.
+In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for English and Chinese academic paper PDFs.
 The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model.
 The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies.
-We also provide a corresponding benchmark named PaperPDF.
-The experimental results substantiate the effectiveness and efficiency of our approach on the task of long multimodal PDF understanding.
+We also provide a corresponding bilingual benchmark named PaperPDF.
+We train PDF-WuKong on PaperPDF dataset, complemented by general-domain document question-answering datasets.
+Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding.
 PDF-WuKong significantly outperforms potential open-source models that may be applied to this task.
@@
   • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient PDF question answering.
-  • We propose a PDF multimodal question answering dataset (PaperPDF) with $1M$ QA pairs for training and $6k$ QA pairs for evaluation.
+  • We propose a bilingual PDF multimodal question answering dataset (PaperPDF) with $1.1M$ QA pairs for training and $10k$ QA pairs for evaluation.
   • Our model significantly surpasses existing open-source models and proprietary products (by an average of $8.6\%$ on F1) on long multimodal PDF understanding.
-Table 1: Comparison of various models for processing multi-page long documents.
 Input modality Type Number of tokens Models
@@
 RAG w/o Linear increase Graph RAG [9], DISC-LawLLM [6], RAPTOR [19]
-Purely visual Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], DocPedia [24], XC2-4KHD [25] and InternVL-V1.5 [11]
-Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [26]
+Pure vision Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], XC2-4KHD [24], InternVL-V1.5 [11]
+Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [25], CREAM [26]
 Text and images Unlimited-page w/o Linear increase PDF-WuKong (Ours)
+Table 1: Comparison of various models for processing multi-page long documents.
 2 Related Works
 2.1 Document Understanding Datasets
-Earlier document understanding datasets only focused on the NLP tasks such as summarization [27], and QA [28] of plain text.
-Meanwhile, there were several visual document datasets mainly aimed at text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34].
-Recently, more datasets have been proposed for multimodal document QA across various scenarios.
-For example, DocVQA [35] and OCRVQA  [36] provide single-page QA data on books and business documents.
-ChartQA [37] and ChartX [38] focus on the visual reasoning for chart documents.
-ArXivQA [39] extracts scientific figure-caption data from the arxiv papers to enhance the academic ability of MLLMs.
-InfoVQA [40] contains many infographic documents which are a combination of textual, graphical and visual elements.
-However, these visual document datasets only define the single-page VQA task, and current MLLMs [41, 11] have achieved remarkable performance on it.
-There are also some multi-page QA datasets [15, 17, 14] that require the model to understand the content relationship via multi-hop reasoning and capture the crucial information from multi-page documents.
-MP-DocVQA [14] extends DocVQA [35] by adding the context pages.
-DUDE [17] constructs multi-page QA data from multi-industry and multi-domain documents.
-Recently, DocGenome [42] was constructed as a scientific document benchmark.
-MM-NIAH [18] is a benchmark evaluating the capability of MLLMs to comprehend long multimodal documents, which requires the model to answer according to the key information scattered throughout the document.
-However, the answers in these datasets lack evidence and cannot provide reliable interpretability, especially for questions that require referring to multiple pieces of evidence in long documents.
+Early datasets focused on NLP tasks like summarization [27] and QA [28] of plain text, while visual document datasets targeted text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34].
+Recent multimodal document QA datasets include DocVQA [35] and OCRVQA [36] for single-page documents, ChartQA [37] and ChartX [38] focusing on visual reasoning in charts.
+Datasets like ArXivQA [39] and InfoVQA [40] enhance MLLMs’ abilities on academic and infographic documents.
+However, these datasets are limited to single-page tasks, and current MLLMs [41, 11] perform well on them.
+Multi-page QA datasets like MP-DocVQA [14], DUDE [17], DocGenome [42], and MM-NIAH [18] require understanding content relationships via multi-hop reasoning.
+Yet, answers in these datasets lack evidence and reliable interpretability, especially for questions needing multiple pieces of evidence from long documents.
 2.2 Document Understanding Methods
-Existing document understanding methods typically focus on either plain text or a limited number of document images.
-Methods that rely on pure text modality aim to process documents by first converting them into plain text through document parsing or Optical Character Recognition (OCR).
-They then employ efficient long-context mechanisms to handle long texts, such as sparse attention [4], memory networks [5], or position interpolation [3].
-Besides, the methods based on retrieval-augmented generation (RAG) [6, 19, 9] also show impressive capabilities for long texts.
-While these approaches can integrate visual elements by including image captions or transforming images into natural language descriptions, they struggle with fine-grained understanding of visual information.
-This restricts their effectiveness in tasks requiring detailed interpretation of textual and visual components within documents.
-Another solution is visual document understanding with purely visual input, treating each document page as an image.
-Many MLLMs such as UniDoc [20], mPLUG-DocOwl [21] and Vary [12] can perform this task in an OCR-free manner.
-Vary [12] employs an extra SAM-style [43] vision vocabulary specific to document and chart data, enabling direct encoding of entire pages with high compression ratios.
-Other researchers have advanced the understanding of high-resolution document pages by dividing input images into smaller patches, such as UReader [22], TextMonkey [10], and LLaVA-NeXT [23].
-DocPedia [24] processes high-resolution images in the frequency domain via the DCT transformation.
-InternLM-XComposer2-4KHD [25] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration to capture more details.
-The reliance on high resolution results in a higher number of tokens and cannot be extended to multi-page documents.
-Besides, there are several models specifically designed for multi-page documents.
-Hi-VT5 [14] and GRAM [16] are two professional models for multi-page QA.
-They combine image tokens of multiple pages through hierarchical transformer architecture [14] or global-local reasoning [16], being able to handle up to 20 pages.
-Fox [13] unifies all image tokens of up to 8 pages into a sequence to achieve multi-page QA.
-mPLUG-DocOwl2 [26] compress each high-resolution document image into 324 tokens, guided by low-resolution global visual features.
-These models encode each page separately and then perform page-level visual feature interactions.
-However, more pages generate more visual tokens, which brings greater resource consumption to LLM.
-This makes them inefficient and unable to process longer documents.
-Given this, we propose first to parse the lengthy document into interleaved text and image content and then perform sparse sampling in an end-to-end manner.
+Existing methods focus on plain text or limited document images.
+Text-based approaches convert documents into plain text using OCR and then employ long-context mechanisms like sparse attention [4], memory networks [5], or position interpolation [3].
+Retrieval-augmented generation methods [6, 19, 9] also handle long texts effectively.
+These approaches struggle with fine-grained visual understanding.
+Another solution, visual document understanding, treats each page as an image.
+MLLMs like UniDoc [20], mPLUG-DocOwl [21], and Vary [12] perform OCR-free understanding.
+Models such as UReader [22] and TextMonkey [10] divide high-resolution pages into patches.
+InternLM-XC2-4KHD [24] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration.
+However, reliance on high resolution increases token counts and isn’t scalable to multi-page documents.
+For multi-page documents, models like Hi-VT5 [14], GRAM [16], Fox [13], CREAM [26] and mPLUG-DocOwl2 [25] encode pages separately and perform page-level interactions.
+More pages generate more visual tokens, increasing resource consumption and inefficiency for longer documents.
+Thus, we propose parsing documents into interleaved text and images, followed by sparse sampling in an end-to-end manner.
 Tab.
-1 summarizes the attributes of each type of method.
+1 summarizes these methods.
 3 Methodology
@@
 3.1 Overview
-In order to achieve multimodal understanding of long PDF documents and alleviate the drawbacks of existing models that treat PDF documents as a single modality of plain text or images, we propose PDF-WuKong.
-The core motivation is that users’ queries are often only related to a small number of text blocks or diagrams in a long document.
-Therefore, to improve the accuracy and efficiency of MLLM, we design a sparse sampler integrated with a multimodal large language model in an end-to-end manner.
-Specifically, our pipeline consists of three parts: a document parser, a sparse sampler and a large language model, as shown in Fig.
+Our pipeline consists of three components: a document parser, a sparse sampler, and a large language model, as shown in Fig.
 2.
-The document parsing stage first converts the input PDF document into machine-readable content of interleaved text and images, Then, the sparse sampler encodes the text blocks and images separately and caches their embeddings.
-When a user inputs a query, the most relevant content can be sampled using a simple similarity measure.
-Finally, the query and the sampled tokens are input into the LLM to generate the answer.
-Algorithm 1 shows the detailed steps of this method.
+The document parsing stage converts input PDFs into machine-readable content with interleaved text and images.
+The sparse sampler then encodes and caches embeddings for text blocks and images separately.
+When receiving a user query, it retrieves the most relevant content through similarity matching.
+Finally, the query and sampled tokens are fed into the LLM for answer generation.
+The detailed procedure is outlined in Algorithm 1 in the appendix.
+3.2 Document Parsing
+Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure.
+By default, text blocks are organized into paragraphs, and all figures and tables are saved as images.
+These text and images are finally reorganized into an XML file in reading order.
+This process can be completed using existing open-source PDF parsing tools.
+During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
+3.3 Sparse Sampling
+For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems.
+The first is the problem of computing efficiency.
+The consumption of computing resources will increase dramatically.
+The second is the problem of inaccurate attention.
+Key information related to the user query is easily submerged by a large amount of irrelevant content.
+It is difficult for the model to accurately locate and extract important information in a huge token sequence.
+Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
+During the training, for the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query.
+Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec.
+4).
+We randomly select two text blocks and two images from the remaining text blocks and images as negative samples.
+Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$.
+An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
+Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query.
+The goal is to enable the model to capture the document content that is most relevant to the query.
+The contrastive learning loss is:
+$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}}{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}+\displaystyle\sum_{e_{j}\in E_{N}}e^{\frac{\text{sim}(e_{q},e_{j})}{\tau}}},$ (1)
+where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples.
+$\tau$ is the temperature parameter that controls the scale of the similarity scores.
+$P$ is the number of positive samples.
+By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart.
+It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
+During the inference, we pre-encode all text blocks and images and cache all candidate embeddings.
+When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings.
+Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query.
+Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences.
+Moreover, the multimodal embedding cache further optimizes inference time.
+The pseudocodes for training and inference are shown in the appendix.
+3.4 Answer Generation
+At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency.
+Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
+Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process.
+Thus, the sparse sampler shares the same vision encoder with the MLLM.
+They can be integrated and trained in an end-to-end manner.
+During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM.
+Besides, the query and instruction are also input into the LLM.
+Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth.
+Finally, the total optimization objective is:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
+PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and QA.
+4 PaperPDF Dataset
+Figure 3: The construction process of PaperPDF based on single evidence and multiple evidence.
+4.1 Overview
+Our motivation for creating PaperPDF is threefold: (1) In long document QA contexts, answers often derive from specific segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multimodal PDF QA dataset.
+Therefore, we introduce a method for automatically generating question-answer pairs from long documents and present PaperPDF, a dataset designed for both training and evaluation.
+4.2 Dataset construction strategy
+The PaperPDF dataset is constructed through a four-step process as shown in Fig.
+3: document parsing, evidence extraction, QA generation, and data filtering.
+First, $100k$ PDFs are parsed to extract text and image chunks.
+Then we extract some evidence from each PDF according to our specific rules.
+Given the evidence and the generation prompt as the input, we use Gemini Pro [43] to generate questions and answers for the training set due to its free and rapid accessibility.
+GPT-4V [44] is used for the test set to ensure high evaluation quality.
+After obtaining the preliminary dataset, we conduct automatic filtering and remove abnormal data.
+For the test data, we further perform manual filtering from multiple aspects to ensure its quality.
+According to different evidence extraction rules, the triplets in PaperPDF can be divided into two categories.
+Single-evidence Q-E-A triplets.
+The questions in these triplets can be answered based on a single text chunk or a diagram.
+Therefore, the evidence can be categorized into Text-only and Image-only.
+These triplets enable PDF-WuKong to initially acquire the capabilities of sparse sampling and long multimodal document understanding.
+Multi-evidence Q-E-A triplets.
+These triplets require reasoning across multiple chunks, involving combinations of text and images.
+The types include Image-text (derived from a text chunk and associated images), Section (generated from all chunks in a section), and Cross-paragraph (involving related paragraphs across a document).
+For the third type, semantic summarizations for each paragraph are conducted first, followed by a selection of several related text chunks for QA generation.
+These multi-evidence triplets enhance our model’s multi-hop reasoning capability.
+In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The dataset statistics are shown in Tab. 2. The appendix contains more data statistics.
+Category Train (En/Zh) Test (En/Zh)
+Single Text-only 249k/12k 2939/296
+Image-only 21K/40k 212/1018
+Multi Image-text 250k/53k 2566/2150
+Section 499k/7k 255/394
+Cross-paragraph 1.2k/0 118/0
+Table 2: The statistics of PaperPDF in English and Chinese.
+5 Experiments
+5.1 Implementation Details
+The LLM and the vision encoder are initialized from IXC2-VL-4KHD [24] and the maximum number of dynamic tiles is set as 16.
+The text encoder is initialized from BGE-M3 [45].
+We train the model by leveraging several document datasets including PaperPDF, DocVQA [35], ChartQA [37], InfoVQA [40], MPDocVQA [46], and DUDE [17].
+Before both training and testing, PaperPDF is parsed using Grobid [47] and MinerU [48], while the other datasets are processed following their default instructions.
+The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5.
+We set the top 5 sampling results as input to the LLM.
+For the model evaluations on the PaperPDF dataset, we use three objective metrics ANLS, F1, and Rouge to report the quantitative results on the full test set.
+To evaluate the semantic correctness of the answer, we introduce GPT-Acc to determine whether the output is correct.
+Considering the expensive cost of GPT-4 evaluation, we randomly selected two subsets with 50 English PDFs and 30 Chinese PDFs for GPT-Acc calculation.
+They contain 488 and 317 QA samples, respectively.
+Model #Param English Chinese Token
+ANLS F1 Rouge GPT-Acc ANLS F1 Rouge GPT-Acc
+Plain Text Solution
+IXC2-VL [49] 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
+IXC2-VL-RAG [49] 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
+InternVL2 [11] 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
+InternVL2-RAG [11] 8B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
+Pure Vision Solution
+Hi-VT5 [46] 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
+IXC2-VL [49] 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
+InternVL2 [11] 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
+Parsed Interleaved Text and Images
+IXC2-VL [49] 8B 27.8 31.2 32.6 37.7 22.5 29.5 29.2 31.4 6217
+InternVL2 [11] 8B 33.4 36.2 36.6 54.3 28.5 40.6 42.0 54.7 6220
+PDF-WuKong (ours) 8.5B 41.9 43.5 40.9 77.5 40.9 47.8 48.6 57.8 2107
+Table 3: Performance comparison with open-source models for long PDF understanding on PaperPDF. The best results are marked bold and the second results are underlined.
+Model ANLS F1 Rouge GPT-Acc
+Gemini pro [50] 26.6 29.0 29.8 67.9
+Kimi [51] 28.5 33.6 31.1 74.7
+ChatGLM [52] 31.2 35.4 32.0 73.5
+Qwen [53] 36.0 40.3 35.5 78.1
+PDF-WuKong (ours) 41.8 43.2 40.7 77.5
+Table 4: Performance comparison with commercial products (tested in Sep 2024) for long PDF understanding.
+The results are tested on a subset of 50 English PDFs.
+Note: These are the products based on the models rather than the models themselves.
+5.2 Long PDF Understanding
+To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset.
+Due to limited capabilities of traditional document understanding models  [46, 16, 54, 55, 56] and the huge resource costs in advanced MLLMs  [49, 24, 11], achieving deep understanding of lengthy PDF documents remains a highly challenging task.
+To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content.
+For some baselines with plain text input modality, we also report results based on retrieval enhancement techniques  [45].
+The experimental results are reported in Tab.
+3 and Tab.
+4.
+Several key conclusions can be drawn from the results.
+Firstly, inputting parsed interleaved text-image content generally outperforms multiple page images.
+The main reason is the limited input resolution and number of tokens that the model can accept.
+Secondly, when handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content.
+It is obvious that diagram information in the PDF plays a crucial role for document comprehension.
+This also indicates that PaperPDF places rigorous requirements on the visual information within documents.
+Additionally, we observe that for the InternVL2 model, the approach of parsing interleaved text-image content as input outperforms InternVL2-RAG.
+However, the opposite conclusion is drawn for IXC2-VL.
+We hypothesize that this discrepancy may be due to the max length setting in IXC2-VL, which could cause the model to overlook some critical information.
+Finally, benefiting from the inclusion of the spare sampler, our proposed PDF-WuKong model not only surpasses the existing state-of-the-art open-source model InternVL2 by approximately 7% on both the Chinese and English subsets, but also demonstrates competitive performance comparable to proprietary products.
+Moreover, due to the integration of the sparse sampler, PDF-WuKong maintains efficiency in inference token cost.
+Model # param ANLS F1 ROUGE
+Qwen-VL [57] 9.6B 26.4 19.6 18.3
+Monkey [58] 9.8B 30.0 24.4 22.3
+mPLUG-Owl2 [25] 8.2B 19.5 20.3 22.7
+Emu2-Chat [59] 37B 26.0 24.4 23.4
+MiniCPM-2.5 [60] 8.5B 31.8 28.2 24.8
+IXC2-VL [49] 8B 23.4 20.8 21.3
+IXC2-4KHD [24] 8B 24.5 20.0 18.0
+CogVLM2 [61] 17B 24.8 27.4 26.3
+[…]
+[…]
+Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on the Single-Evidence Subset.
+$\dagger$ denotes the page image input, aligning with other models in the table; * indicates that our model utilizes parsed content as input.
+To compare with more open-source document MLLMs, considering that most of these models can only handle single-page documents, we construct a subset of the PaperPDF benchmark, only containing test samples with single evidence.
+Therefore, we provided all models with only one page containing the evidence as their input.
+The pages are input in the form of images.
+Our PDF-WuKong can accept input in two formats.
+One is the page image and another is the parsed page.
+As shown in Tab.
+5, our model’s capability on this subset is significantly better than other document models.
+Moreover, the document parsing-based paradigm used in our approach is superior to the purely visual paradigm.
+5.3 Document-oriented VQA
+To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks and compare PDF-WuKong with other representative models.
+First, we evaluate the performance of PDF-WuKong on single-page document datasets [35, 37, 40].
+As shown in Tab.
+6, our model achieves comparable performance on single-page document understanding.
+This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
+In addition, we assess the performance of traditional specialized models and MLLMs on two existing multi-page document QA datasets.
+The results shown in Tab.
+7 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 [25].
+Notably, on complex multi-page document datasets like DUDE [17], PDF-WuKong outperforms GPT-4V [44].
+This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
+Furthermore, we conduct zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18].
+As shown in Tab.
+8, our model uses the fewest parameters yet achieves the second-best performance.
+Although InternVL-V1-5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model.
+Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases.
+At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
+Model DocVQA ChartQA InfoVQA
+Qwen-VL [57] 65.1 65.7 35.4
+Monkey [58] 66.5 65.1 36.1
+Text-Monkey [10] 73.0 66.9 28.6
+MiniCPM-V-2.5 [60] 84.8 - -
+Vary-base [12] 76.3 66.1 -
+TextHawk [62] 76.4 66.6 50.6
+IXC2-4KHD-16 [24] 84.9 80.1 60.8
+DocOwl 2 [25] 80.7 70.0 46.4
+CREAM [26] 79.4 - 53.6
+PDF-WuKong (ours) 85.1 80.0 61.3
+Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks.
+Model MP-DocVQA DUDE
+LayoutLMv3 [54] 55.1 20.3
+Longformer [55] 55.1 27.1
+BigBird [56] 58.5 26.3
+Hi-VT5 [46] 61.8 35.7
+DocFormerv2 [63] 76.4 48.4
+GRAM [16] 83.0 53.4
+GPT-4V (2024-06) [44] - 53.9
+Idefics3-8B [64] 67.2 38.7
+DocOwl2 [25] 69.4 46.7
+CREAM [26] 65.3 52.5
+PDF-WuKong (ours) 76.9 56.1
+Table 7: Performance comparison with other DocVLMs for multi-page document understanding.
+Model #param Overall 1K 4K 16K 64K
+Emu2-Chat [59] 37B 8.8 38.9 18.2 0.0 0.0
+VILA1.0-13b [65] 13B 15.7 41.9 33.2 8.6 0.1
+llava-v1.6-13b [23] 13B 16.9 43.7 34.9 13.6 0.0
+llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
+InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
+InternVL1.5-RAG [11] 45B 46.1 59.5 50.1 44.9 39.3
+PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
+Table 8: Performance comparison with other DocVLMs on MM-NIAH. The evaluation approach aligns with the benchmark.
+5.4 Ablation Study
+To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies.
+These experiments are based on English PaperPDF for training and evaluation to avoid interference from other factors.
+Sparse sampler
+To assess the effectiveness of the sparse sampler, we compared models trained with and without it.
+Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the large amount of irrelevant information.
+Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab.
+9, by efficiently selecting the most relevant content for each query.
+Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately.
+This indicates that our end-to-end optimization of multimodal representation and question answering can further promote the document understanding ability of MLLM.
+Sparse Sampler End-to-End ANLS F1 ROUGE
+11.1 5.1 5.0
+40.3 42.3 39.8
+42.6 43.6 40.2
+Table 9: Ablation study on the impact of sparse sampler
+Dataset
+We retrain PDF-WuKong on various subsets of our English PaperPDF, and the results are shown in Tab.
+10.
+Increasing the amount of training data leads to consistent improvements in the model’s accuracy, proving that our dataset follows scaling laws.
+In addition, we verify the effectiveness of our two data construction methods.
+Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
+Dataset ANLS F1 ROUGE
+100 k 38.7 40.1 37.5
+500 k 41.6 43.5 40.8
+1 M 42.6 43.6 40.2
+Single (200k) 39.2 40.4 38.1
+Multi (100k) + Single (100k) 40.0 41.6 38.9
+Table 10: Ablation study on dataset setting
+Document length
+To understand the impact of document length on model performance and efficiency, we divide the test set into subsets based on the number of pages per document.
+Results in Fig.
+4 demonstrate that our model’s performance and token counts remain relatively stable across documents of varying lengths.
+This stability indicates that the sparse sampler effectively reduces the input size to a reasonable level, regardless of the original document length.
+In contrast, the baseline MLLM without the sparse sampler is unable to handle long documents effectively.
+Its performance deteriorates significantly as the document length increases.
+The number of tokens also increased dramatically, resulting in huge resource consumption.
+These findings highlight the robustness of our model in processing long documents without sacrificing accuracy or incurring extra computational costs.
+Figure 4: Ablation study of different document length
+Sampling strategy
+We explore the impact of different numbers of text blocks or diagrams selected by the sparse sampler.
+As shown in Tab.
+11, setting a small top $k$ can lead to missing crucial information needed for accurately answering queries, thus reducing performance.
+Conversely, a larger top $k$ introduces redundant information and increases computational costs without significantly enhancing accuracy.
+To strike a balance between performance and resource efficiency, we use the top 5 as the default setting.
+Sampling chunks ANLS F1 ROUGE Tokens
+Top 1 39.20 38.28 35.26 1186
+Top 3 42.09 43.06 39.69 1452
+Top 5 42.59 43.63 40.19 1789
+Top 10 43.01 44.22 40.67 2386
+Top 15 43.19 44.57 42.08 2704
+Top 20 43.42 45.02 42.30 3364
+Table 11: Ablation study of different sampling strategy
+6 Conclusion
+We have presented PDF-WuKong, a novel MLLM that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images.
+By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length.
+We also constructed PaperPDF, a bilingual dataset with $1.1M$ question-answer pairs for training and $10k$ pairs for evaluation, specifically tailored for academic PDFs.
+Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on the long multimodal PDF understanding task.
+Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
+Appendix A Algorithm
+Algorithm 1 shows the detailed inference process of PDF-WuKong.
+The training pipeline is shown in Algorithm 2.
+Our PDF-WuKong can achieve efficient and accurate understanding of long PDFs with end-to-end sparse sampling.
 Algorithm 1 Inference pipeline for PDF-WuKong
@@
 13:Return the generated answer $a$.
+Appendix B Visualization Results
+B.1 Qualitative Comparison with Other Products
+The qualitative comparison between PDF-WuKong and other proprietary products on the long PDF understanding task is presented in Fig.
+5.
+PDF-WuKong achieves significant performance advantages over others such as Qwen [66], ChatGLM [67], Kimi [68], and Gemini-pro [69].
+The success stems from its ability to effectively sample key evidence relevant to the question and its strong multimodal understanding capabilities.
+This observation further validates the effectiveness of our method in enhancing the long PDF understanding capabilities of MLLMs.
 Algorithm 2 Training pipeline for PDF-WuKong
@@
 14:Return the final loss $\mathcal{L}_{\text{total}}$.
-3.2 Document Parsing
-Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure.
-By default, text blocks are organized into paragraphs, and all figures and tables are saved as images.
-These text and images are finally reorganized into an XML file in reading order.
-This process can be completed using existing open-source PDF parsing tools.
-During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
-3.3 Sparse Sampling
-For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems.
-The first is the problem of computing efficiency.
-The consumption of computing resources will increase dramatically.
-The second is the problem of inaccurate attention.
-Key information related to the user query is easily submerged by a large amount of irrelevant content.
-It is difficult for the model to accurately locate and extract important information in a huge token sequence.
-Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
-For the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query.
-Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec.
-4).
-We randomly select two text blocks and two images from the remaining text blocks and images as negative samples.
-Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$.
-An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
-Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query.
-The goal is to enable the model to capture the document content that is most relevant to the query.
-The contrastive learning loss is:
-$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{{\text{sim}(e_{q},e_{i})}/\tau}}{e^{{\text{sim}(e_{q},e_{i})}/\tau}+\displaystyle\sum_{e_{j}\in E_{N}}e^{{\text{sim}(e_{q},e_{j})}/\tau}},$ (1)
-where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples, respectively.
-$\tau$ is the temperature parameter that controls the scale of the similarity scores.
-$P$ is the number of positive samples.
-By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart.
-We align the feature space of the text encoder to the pre-trained vision encoder with frozen parameters.
-It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
-During the inference, we pre-encode all text blocks and images and cache all candidate vector embeddings.
-When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings.
-Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query.
-Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences.
-Moreover, the multimodal embedding cache further optimizes inference time.
-3.4 Answer Generation
-At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency.
-Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
-The inference process is shown in Algorithm 1.
-Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process.
-Thus, the sparse sampler shares the same vision encoder with the MLLM.
-They can be integrated and trained in an end-to-end manner.
-During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM.
-Besides, the query and instruction are also input into the LLM.
-Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth.
-Finally, the total optimization objective is:
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
-PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and question-answering. The training pipeline of PDF-WuKong is shown in Algorithm 2.
-4 PaperPDF Construction
-Figure 3: The construction process of PaperPDF based on single evidence and multiple evidence.
-4.1 Overview
-The reasons for constructing our dataset include the following two points.
-In most long document Q&A scenarios, the basis for an answer to a given question is typically composed of a single element or a combination of multiple elements, while the remaining information may act as noise, interfering with the reasoning of MLLMs.
-Additionally, existing document question-answering datasets are either limited to single-page formats or provide only page-level ground truth, which hinders the training of the sparse sampler in PDF-WuKong.
-Consequently, we propose a reliable method for generating high-quality question-answer pairs of long document and develop the dataset PaperPDF for training and evaluation based on this method.
-Category Training Test
-Single Text-only 249k 2939
-Image-only 21K 212
-Multi Image-text 250k 2566
-Section 499k 255
-Cross-paragraph 1.2k 118
-Table 2: The statistic of PaperPDF.
-Text-only and Image-only indicate that the QA pairs are generated based on either a single text paragraph or an image extracted from the PDF.
-Image-text, Section, and Cross-paragraph denote that the QA pairs are generated from a paragraph and its corresponding references, an entire section, or non-contiguous paragraphs, respectively.
-4.2 Detail of Construction
-The construction process of PaperPDF are presented in Fig.
-3.
-It can be divided into four steps: structured parsing, rule-based extraction, prompt construction, and filtering.
-We obtained 89k PDF documents from the arXiv repository as our document set $\mathbb{D}$.
-For each document $D$ in $\mathbb{D}$, we first employed Grobid to parse the document and extract its text chunks $\{T_{1},T_{2},\ldots,T_{m}\}$ and images chunks $\{I_{1},I_{2},\ldots,I_{n}\}$.
-Subsequently, predefined rules are used to randomly select chunks from the document, which may consist of text chunks $T_{i}$, chart chunks $I_{j}$ or a combination of both.
-The selected chunks will be input into commercial MLLM products according to different prompt templates and then, the question $Q$ related to the input chunks, along with the corresponding answer $A$, is subsequently generated.
-Notably, For the training set, we used Gemini for generation due to its free and rapid accessibility, while the high-performance GPT-4V was employed to construct the test set, ensuring the validity and robustness of the evaluation.
-Finally, we devise a set of rules to filter the generated training and testing sets automatically.
-The removing rules for the samples include too-short questions, too-long answers, non-English text, etc.
-Further manual checking is conducted on test set to ensure the reliability of the evaluation.
-PaperPDF consists of two types of QA pairs: Single-evidence and Multi-evidence.
-Single-evidence QA pair means that the question can be answered based on a single text chunk or image chunk in the long document.
-It can be systematically categorized into two types: text-only and image-only.
-In text-only generation, a text chunk is input into MLLM, whereas in image-only generation, an image chunk is utilized.
-The generation process for Single-evidence QA pairs is relatively straightforward and cost-effective, primarily aimed at training the foundational multimodal understanding and question-answering capabilities of MLLMs.
-Multi-evidence QA pairs.
-Unlike single-evidence QA pairs, Multi-evidence QA data exhibits greater complexity, both in terms of characteristics and generation process.
-The answers to such questions typically rely on multiple text chunks, image chunks, or any combination thereof in document.
-Multi-evidence QA pairs consist of Image-text QA pairs, Section QA pairs, and Cross-paragraph QA pairs.
-The Image-text QA pairs are generated based on a paragraph and its corresponding image references.
-These pairs require consideration of the interrelationship between the text and the image to ensure the answers capture multimodal information.
-Section QA pairs are designed to train the MLLM in integrating and understanding information within a section and are generated from all chunks within a section.
-Cross-paragraph QA pairs involve the most complex generation process.
-First, the entire document is input into the MLLM for paragraph-level semantic summarization, followed by the selection of semantically related text chunks.
-Finally, multiple related chunks are randomly selected and re-input into the MLLM to generate the final QA pairs.
-This type of data primarily focuses on training the sampler’s performance, enhancing its sparse sampling capability.
-Although the generation of Multi-evidence QA pairs is relatively complex and may require multiple MLLM calls, their presence significantly improves the performance of the sampler and strengthens the MLLM’s ability to comprehend cross-chunk information.
-Totally, we generated 1.5M samples for training and 8K samples for testing.
-After filtering, the final dataset consists of 1M training data points and 6K test data points for subsequent training and evaluation.
-The statistics of PaperPDF are presented in Tab.
-2.
-Figure 4: Examples from PDF-WuKong and other proprietary products
-5 Experiments
-[…]
-Model #param ANLS F1 ROUGE Token Latency GPT Acc
-[…]
-[…]
-[…]
-[…]
-[…]
-Closed Qwen [46] - 36.0 40.3 35.5 - - 78.1
-Kimi [47] - 28.5 33.6 31.1 - - 74.7
-Gemini pro [48] - 26.6 29.0 29.8 - - 67.9
-ChatGLM [49] - 31.2 35.4 32.0 - - 73.5
-[…]
-5.1 Implementation Details
-[…]
-Table 4: Description of the input format used in the experiment
-Symbol Description
-* Input the parsed content of the pdf file.
-[…]
-$\dagger$ Input the entire page image.
-$\ddagger$ Input the entire page image and OCR content.
-5.2 Long PDF Understanding
-To assess the effectiveness of our model in understanding long PDF documents, we conducted comprehensive experiments comparing it with both open-source models and commercial products on PaperPDF dataset.
-Due to the limited availability of open-source models capable of handling this task, we firsly evaluate two such models that can process multi-page PDF documents.
-Considering that IXC2-VL [45] cannot understand such long image-text interleaved sequences, we also report the results with the input of pure OCR content.
-To further demonstrate the advantages of our multi-modal sampler, we demonstrate the evaluation of IXC2-VL with RAG where we feed the top 5 text paragraphs retrived by BGE-M3 [50].
-The results on the PaperPDF dataset are reported in Tab.
-3.
-The findings indicate that our model significantly outperforms existing open-source models across various evaluation metrics.
-Additionally, thanks to the introduction of the sparse sampler, the number of tokens that our model’s MLLM needs to process is substantially fewer than that of other models.
-Moreover, we compared our model with several commercial products that allow users to input PDF documents and questions via web interfaces to generate answers.
-Considering the cost of manual testing, we randomly selected 50 PDFs from the PaperPDF dataset for evaluation, and the results are presented in Tab.
-3.
-Our model clearly outperforms mainstream PDF question-answering products.
+Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on.
+B.2 Visualization Results on Chinese Documents
+In addition to the outstanding performance on English documents, PDF-WuKong also demonstrates remarkable capabilities on Chinese documents.
 Fig.
-4 shows some examples from PDF-WuKong and other proprietary products.
-Our model can accurately retrieve the evidence needed to answer user questions and generate accurate answers based on relevant text segments or diagrams.
-Since these commercial solutions do not provide detailed code or technical reports, we could not obtain information about their parameter sizes or token counts, and uploading documents directly to their websites does not allow us to accurately assess the models’ latency.
-Finally, we evaluated the GPT accuracy of all open-source and proprietary models on this subset benchmark.
-The results show that our model has significant advantages in both textual similarity and semantic understanding.
-To compare with more open-source large document models, considering that most of these models can only handle single-page documents, we constructed a subset of the PaperPDF benchmark containing only test samples with single evidence.
-Therefore, we provided all models with only the page containing the evidence as input; note that the pages are input in the form of images.
-These models and our PDF-WuKong have not been trained on such data.
-In this setting, our sparse sampler sampled all the content of that page.
-As shown in Tab.5, our model’s zero-shot capability on this subset is significantly better than other document large models.
-Table 5: Performance comparison with other DocVLMs for PDF multimodal understanding on Single-Evidence Subset of PaperPDF.
-($\dagger$) indicates the use of a special input approach.
-For details, refer to Table 4.
-Model # param ANLS F1 ROUGE
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-5.3 Document-oriented VQA
-To validate the generalization capability of our model to other document understanding scenarios, we conducted experiments on several benchmark datasets and compared PDF-WuKong with other representative models.
-First, we evaluated the performance of PDF-WuKong and other open-source models on DocVQA [35], ChartQA [37], and InfoVQA [40], which are all single-page document datasets.
-As shown in Tab.
-6, our model achieved leading performance compared to other open-source models.
-This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question answering tasks.
-In addition, we assessed the performance of traditional specialized models and large-scale models on two existing multi-page document QA datasets.
-The experimental results, presented in Tab.
-7, indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document large model, DocOwl2 [26].
-Notably, on complex multi-page document datasets like DUDE [58], PDF-WuKong outperforms GPT-4V [59].
-This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
-Furthermore, we conducted zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18].
-As shown in Tab.
-8, our model uses the fewest parameters yet achieves the second-best performance.
-Although InternVL-V1-5-RAG [18] surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model.
-Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, unlike other models that experience significant declines.
-At a context length of 64K, PDF-WuKong achieves the best results, demonstrating its robustness in handling long-context multimodal inputs.
-Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
-Model Doc. Chart. Info.
-Closed Gemini Pro [60] 88.1 74.1 75.2
-GPT-4V [61] 88.4 78.5 75.1
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 7: Performance comparison with other DocVLMs for multi-page document understanding. ($\dagger$) indicates the use of a special input approach. For details, refer to Table 4.
-Model MP-DocVQA DUDE
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 8: Performance comparison with other DocVLMs on MM-NIAH. The input approach aligns with the benchmark.
-Model #param Overall 1K 4K 16K 64K
-Emu2-Chat [55] 37B 8.8 38.9 18.2 0.0 0.0
-VILA1.0-13b [69] 13B 15.7 41.9 33.2 8.6 0.1
-llava-v1.6-13b [70] 13B 16.9 43.7 34.9 13.6 0.0
-llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
-InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
-InternVL1.5-RAG [18] 45B 46.1 59.5 50.1 44.9 39.3
-PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
-5.4 Ablation Study
-To comprehensively evaluate the effectiveness of our proposed model components, we conducted ablation studies focusing on datasets, the impact of the sparse sampler, sampling strategies, and document length.
-Below, we present the findings from each of these experiments.
-Datasets
-We conducted experiments to evaluate the impact of training data size and composition on model performance.
-As shown in Tab.
-9, increasing the amount of training data led to consistent improvements in the model’s accuracy on the PaperPDF benchmark.
-Specifically, the model trained with larger datasets demonstrated better comprehension and answer accuracy.
-Moreover, under the same data volume, training sets that included multi-evidence annotations outperformed those containing only single-evidence.
-This suggests that exposure to diverse and complex evidence during training enhances the model’s ability to handle queries requiring multiple pieces of information, thereby improving overall performance.
-Table 9: Ablation study on dataset setting
-Dataset ANLS F1 ROUGE
-100 k 38.7 40.1 37.5
-500 k 41.6 43.5 40.8
-1 M 42.6 43.6 40.2
-Impact of sparse sampler
-To assess the effectiveness of the sparse sampler, we compared models trained with and without it.
-Without the sparse sampler, the MLLM struggled to process long documents with interleaved text and images, resulting in poor performance due to the overwhelming amount of irrelevant information.
-Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab.
-10, by efficiently selecting the most relevant content for each query.
-Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately.
-This indicates that our integrated design allows the two components to mutually enhance each other, optimizing both the retrieval of pertinent information and the generation of accurate answers.
-Table 10: Ablation study on the impact of sparse sampler
-Sparse Sampler End-to-End ANLS F1 ROUGE
-11.1 5.1 5.0
-40.3 42.3 39.8
-42.6 43.6 40.2
-Sampling strategy
-We explored different sampling strategies by varying the number of top N text paragraphs or figures selected by the sparse sampler.
-As shown in Tab.
-11, setting N too low resulted in missing critical information necessary for accurately answering queries, thereby reducing performance.
-Conversely, a larger N introduced redundant information and increased computational and time costs without significant improvements in accuracy.
-This suggests that there is an optimal range for N that balances the inclusion of essential information with efficiency.
-Selecting the appropriate number of samples is crucial for maximizing performance while minimizing resource consumption.
-Table 11: Ablation study of different sampling strategy
-Sampling chunks ANLS F1 ROUGE Tokens
-Top 1 39.20 38.28 35.26 1186
-Top 3 42.09 43.06 39.69 1452
-Top 5 42.59 43.63 40.19 1789
-Top 10 43.01 44.22 40.67 2386
-Top 15 43.19 44.57 42.08 2704
-Top 20 43.42 45.02 42.30 3364
-Figure 5: Ablation study of different document length
-Document length
-To understand the impact of document length on model performance and efficiency, we divided the test set into subsets based on the number of pages per document.
-Results in Fig.
-5 demonstrate that our model’s performance and time efficiency remained relatively stable across documents of varying lengths.
-This stability indicates that the sparse sampler effectively reduces the input size to a manageable level, regardless of the original document length.
-In contrast, MLLMs without the sparse sampler were unable to handle long documents effectively; their performance and time efficiency deteriorated significantly as the document length increased.
-These findings highlight the robustness of our approach in processing long documents without sacrificing accuracy or incurring additional computational costs.
-6 Conclusion
-We have presented PDF-WuKong, a novel Multimodal Large Language Model that effectively addresses the challenges of understanding long PDF documents containing interleaved text and images.
-By introducing an end-to-end sparse sampling mechanism, our model efficiently extracts the most relevant paragraphs and diagrams in response to user queries, significantly reducing input token size and making the process independent of document length.
-We also constructed PaperPDF, a comprehensive dataset with 1 million question-answer pairs for training and 6,000 pairs for evaluation, specifically tailored for academic PDFs.
-Experimental results demonstrate that PDF-WuKong not only outperforms existing open-source models but also surpasses proprietary products by an average of 8.6% in F1 score on long multimodal PDF understanding tasks.
-Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
-Appendix A Dataset Examples
-Text-only Data Example in the Training Set Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model?
+6 shows its qualitative performance on the Chinese long PDF document understanding task.
+It effectively samples relevant evidence within the PDF and demonstrates the ability to derive accurate answers based on the sampled evidence.
+Appendix C More Dataset Details
+The Q-E-A triplets in the PaperPDF dataset are derived from approximately 60k English and 10k Chinese documents, encompassing nearly 70 disciplines such as computer science, engineering, and materials science.
+To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis.
+The key statistics of the source documents in PaperPDF are presented in Tab.
+12.
+Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents but also features more pages and OCR tokens.
+More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images.
+Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
+For the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence, Fig.
+7 - Fig.
+11 show their prompt engineering and a corresponding example, respectively.
+Figure 6: Examples of PDF-WuKong on Chinese documents. The red box indicates the evidence that the correct answer depends on.
+MPdataset Document Language Page Image Token
+DUDE [17] 5k En 6 - 1831
+MP-Docvqa [46] 6k En 8 - 2026
+PaperPDF (En) 60k En 25 12 11371
+PaperPDF (Zh) 10k Zh 11 7 3413
+PaperPDF 70k En+Zh 23 11 10234
+Table 12: Statistics of the source documents, where Page, Image, and Token represent the average number of pages per document, the average number of images in the document, and the average number of OCR tokens per document, respectively.
+Text-only Question Generation Prompt Task Definition: Create 2 academic questions from a given research paper paragraph.
+Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions.
+Focus on the paragraph,not the entire paper.
+If the paragraph lacks valid information,return ‘quit’.
+You should use English.
+Develop 2 questions that: Are no more than 30 words.
+Incorporate knowledge from the paragraph.
+Are answered by text instead of one of the multiple choices.
+Are elicit detailed responses supported by the text.
+Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
+Text-only Answer Generation Prompt Task Definition: Answer a question based on the material given.
+Requirements: The answers should: Be comprehensive and cover all relevant aspects.
+Accurately reflect the paragraph’s information and insights.
+You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than 20 words.
+[thinking procedure]: [A2]: answer2 here, no more than one sentence.
+Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model?
 Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig.
@@
 If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
-Text-image Data Example in the Training Set Query: What is the best performing method for both detection and classification, according to the provided figure?
+Figure 7: Text-only Q-E-A triplets generation prompt and data example.
+Image-only Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided figure or table from a research paper.
+Requirements: The questions must directly reference and integrate information presented in the image and its caption, ensuring a cohesive understanding of the content depicted.
+You should use English.
+Develop 2 questions that are: No more than 30 words.
+Specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables.
+Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’.
+Expected Output: [Q1]:[Q2]:
+Image-only Answer Generation Prompt Task Definition: Answer a question based on an image and its caption from a research paper.
+Requirements: The answers should: Always use English.
+Not infer or speculate on details not explicitly explained by the figures/tables.
+You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words.
+[thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
+Image-only Data Example Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period?
+Figure: Figure 4Refer to caption Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour.
+A benchmark is counted for multiple techniques if two techniques are tied for the lead.
+Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4.
+Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
+Figure 8: Image-only Q-E-A triplets generation prompt and data example.
+Text-image Question Generation Prompt Task Definition: Formulate 2 academic questions based on a provided paragraph, a figure or table from a research paper.
+Requirements: The questions should not rely on the accompanying text but only the figure / table.
+However, you may use the provided text to understand the figure / table.
+You should use English.
+Develop 2 questions that are: No more than 30 words.
+Based on a provided figure of a research paper, without relying on accompanying text for the questions.
+However, you may use the provided text to understand the figure.
+Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’.
+Expected Output: [Q1]:[Q2]:
+Text-image Answer Generation Prompt Task Definition: Answer a question based on the material given.
+Requirements: The answers should be: Always using English.
+Comprehensive and cover all relevant aspects, as presented in the provided text and figure.
+Accurately reflect the information and insights offered by the research paper.
+You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words.
+[thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
+Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure?
+Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1).
+Figure: Table 1Refer to caption Caption: Experimental results of ablation analysis, … Answer 1: Ours.
+Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks.
+This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
+Figure 9: Text-image Q-E-A triplets generation prompt and data example.
+Section Question Generation Prompt Task Definition: Formulate 2 academic questions based on a section from a research paper.
+Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions.
+You should Always use English.
+Develop 2 questions that are: No more than 30 words.
+Requiring an integration of information from all paragraphs and figures/tables in the section.
+Not mentioning the label of the figure/table directly or use words like ‘from the figure/table’.
+Not based on common knowledge or assumptions not supported by the figures and tables.
+Expected Output: [Q1]:[Q2]:
+Section Answer Generation Prompt Task Definition: Answer a question based on the material given.
+Requirements: The answers should be: Always using English.
+Not inferring or speculating on details not explicitly explained by the figures/tables.
+You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: …  [A1]: answer1 here, no more than 20 words.
+[thinking procedure]: …  [A2]: answer2 here, no more than one sentence.
+Section Data Example Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries?
+Text:
+Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers…
+Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method… The images clearly show a contrast of content… Figure: Figure 3Refer to caption Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right).
+Answer 1: Visual contrast in word frequency highlights content focus differences.
+Answer 2: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually.
+Figure 10: Section Q-E-A triplets generation prompt and data example.
+Cross-paragraph Question Generation Prompt Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question.
+This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper.
+Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided).
+Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper.
+Expected Output: [Q]: [Your generated question based on the shared themes or information]
+Cross-paragraph Answer Generation Prompt Task Definition: Given some selected paragraphs from a research paper, and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs.
+Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material.
+Are comprehensive and cover all relevant aspects, as presented in the provided figure.
+Expected Output: [A1]: [Insert a concise answer here, no more than 20 words.] [A2]: [Insert a detailed answer here,including a detailed “thought chain” or reasoning process.]
+Cross-paragraph Data Example Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms?
+Text:
+Paragraph 7:
+Figure 1: On the right: the current path for selecting some algorithm(s), …
+Paragraph 17: Section 2 describes the structure of the proposed method including its parameters and operators.
+Paragraph 20: In this section, the details of an algorithm that generates data for which a given function, … Figure: Figure 1Refer to caption Caption: Sample number of rows and columns 2.
+Sample columns and fill/trim values as needed.
+Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection.
+Answer 2: We can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use.
+This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy.
+Figure 11: Cross-paragraph Q-E-A triplets generation prompt and data example.
+Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model?
+Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig.
+5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig.
+5b).
+Answer 1: It can lead to an unfair advantage for the model.
+Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset.
+If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
+Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure?
 Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1).
@@
 This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
-Section Data Example in the Training Set Query: What kind of information does the factor R i, j hold? Text:
-RECOMMENDATION MODEL
-In this section, we present our Joint Topic-Semantic-aware Matrix Factorization (JTS-MF) model for online social votings, in which social relationship, group a liation, and topic-semantic similarities are combined and taken into account for voting recommendation in a comprehensive manner.Motivated by Locally Linear Embedding [23] which tries to preserve the local linear dependency among inputs in the low-dimensional embedding space, we expect to keep inter-user and inter-voting topic-semantic similarities in latent feature space as well.To this end, in JTS-MF model, while the rating R i, j is factorized as user latent feature Q i and voting latent feature P j , we deliberately enforce Q i and P j to be dependent on their social-topic-semantic similar counterparts, respectively.e graphic model of JTS-MF model is as shown in Figure 5.
-Figure: Figure 5[Uncaptioned image] Caption: Fig.
-5: Graphic Model of JTS-MF.
-Answer 1: The factor R i, j represents the rating that user i gave to voting j.
-Answer 2: The node labeled R i, j is a factor that represents the rating given by user i to a voting j.
-It is the core of the model, representing the essence of user interaction with the voting system.
-This node’s existence suggests that the model is primarily concerned with understanding and predicting user ratings for voting content.
-Text-only Data Example in the Test Set Query: How does the distribution of output image transform tasks among N workers impact the efficiency of the final synchronization task?
-Text: The number of output image transform tasks equals the number of output images times the batch size.
-The tasks are executed by all N workers, such that each worker picks up an arbitrary task and executes it.
-The last output image transform task to finish also executes the final synchronization task, which frees the memory required for the output image transforms.
-Answer 1: The distribution allows for parallel processing, potentially speeding up the final synchronization task.
-Answer 2: Ensures faster task completion, synchronization efficiency relies on all tasks’ completion.
-Text-image Data Example in the Test Set Query: What distinguishing features can be observed between ‘hit’ and ‘miss’ samples for real fingerprints in Fig.
-5 across different benchmarks?
-Text: Images found in these benchmarks can be observed in Fig.
-5 of Section V.As we can see, variability exists not only across modalities, but also within modalities.Moreover, it is rather unclear what features might discriminate real from spoofed images, which suggests that the use of a methodology able to use data to its maximum advantage might be a promising idea to tackle such set of problems in a principled way.
-Figure: Figure 5[Uncaptioned image] Caption: Fig.
-5.
-Examples of hit and missed testing samples lying closest to the real-fake decision boundary of each benchmark.
-A magnified visual inspection on these images may suggest some properties of the problem to which the learned representations are sensitive.
-Answer 1: Hit’ samples for real fingerprints show clearer ridge patterns and quality compared to ‘miss’ samples throughout the benchmarks.
-Answer 2: Hit real fingerprints are clearer, more distinct; Miss real samples are blurred, less defined.
-Only-image Data Example in the Test Set Query: Based on Figure 4, which fuzzing technique consistently leads in coverage across all benchmarks over the 24-hour period?
-Figure: Figure 4[Uncaptioned image] Caption: Figure 4: Number of benchmarks on which each technique has the lead in coverage at each hour.
-A benchmark is counted for multiple techniques if two techniques are tied for the lead.
-Answer 1: FairFuzz consistently leads in coverage across all benchmarks over the 24-hour period in Figure 4.
-Answer 2: FairFuzz, highest coverage benchmark count over 24 hours.
-Section Data Example in the Test Set Query: How does the qualitative evaluation of extractive summarizers using word clouds elucidate the differences in content focus between the original documents and the summaries?
-Text:
-Here we use word cloud representations to give an intuitive interpretation of the content in the generated extractive summarizers.We create word clouds for the two best methods in Section 5.3.In this paper, we used an online tool called WordItOut5 to generate the word cloud representations.In all word clouds presented in this paper, a filter is used to display only the words with minimum frequency of 2.
-Figure 3 shows a word cloud made by the aggregation of all the summaries generated by the PKUSUMSUM-Centroid method.This gives a sense of the content in those summaries.For contrast, we also generate a word cloud for the original news articles without the content of the generated summaries.Specifically, common words are first removed completely and then the word clouds are built with frequencies of surviving words.In essence, this shows what information remains apart from the generated summaries.
-The images clearly show a contrast of content.The summary wordcloud shows “earthquake” as its most prominent word.The image of the articles show less focus.If viewed alone, the reader would not quickly infer the gist of the original content.Similarly, Figure 4 represents “Lead” method.And here we see an even more stark difference.6.
-Comparing Twitter and News Media information.
-Figure: Figure 3[Uncaptioned image] Caption: Figure 3: The word clouds representing summaries generated by PKUSUMSUM-Centroid method (left) and original documents without the content of those summaries (right).
-Answer 1: Word clouds highlight the prominent themes in summaries versus original texts by displaying relative word frequencies visually.
-Answer 2: Visual contrast in word frequency highlights content focus differences.
-Cross-paragraph Data Example in the Test Set Query: How can we leverage the proposed EDO approach to optimize the selection of datasets for specific algorithms, thereby enhancing the overall performance and validity of the algorithms?
-Text:
-Paragraph 7:
-Figure 1: On the right: the current path for selecting some algorithm(s) based on their validity and performance for a given dataset.On the left: the proposed flip to better understand the space in which ‘good’ datasets exist for an algorithm.
-Paragraph 17:
-Section 2 describes the structure of the proposed method including its parameters and operators.
-Paragraph 20:
-In this section, the details of an algorithm that generates data for which a given function, or (equivalently) algorithm, is well-suited is described.This algorithm is to be referred to as “Evolutionary Dataset Optimisation” (EDO).
-Figure: Figure 1[Uncaptioned image] Caption: Sample number of rows and columns 2.
-Sample columns and fill/trim values as needed.
-Answer 1: The EDO approach identifies optimal datasets, improving algorithm performance and validity through tailored data selection.
-Answer 2: By employing the EDO approach, we can systematically evaluate and select datasets that align closely with the specific requirements of the algorithms in use.
-This method analyzes various dataset characteristics, ensuring that the chosen data not only matches the algorithm’s operational parameters but also enhances its predictive accuracy.
-As a result, the performance metrics of the algorithms improve significantly, leading to more valid outcomes and reliable interpretations.
-The focus on tailored dataset selection directly contributes to optimal algorithm efficiency, thereby validating the effectiveness of the EDO approach in practical applications.
-Appendix B Prompt Engineering
-Text-only QA Generation Prompt for the Training Set Task Definition: Generate 3 high quality academic questions and corresponding answers based on the provided text/image/table.
-Requirements: The questions should require very complex reasoning and global understanding and the answer should be detailed and can be answered by the material, do not add extra information.
-Remember it is only a paragraph rather than a whole passage so do not ask global question about the study.
-Generate question-answer pairs instead of multiple-choice questions.
-Do not generate global question about the paper such as main idea or abstract.
-You should use English.
-Expected Output: List your Q and A as: [Q1]: [A1]: [Q2]: [A2]: [Q3]: [A3]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea.
-Figure 6: Text-only QA generation prompt for the training set.
-Text-image QA Generation Prompt for the Training Set Task Definition: Generate 2 high-quality academic questions and corresponding answers based solely on a provided figure of a research paper, without relying on accompanying text for the questions.
-However, you may use the provided text to understand the figure.
-Ensure these questions demand complex reasoning, focusing exclusively on details within the figure that are critical for understanding.
-The answers must be detailed, directly reflecting the content in the figure and capable of being derived solely from the figure without extraneous information.
-Requirements: Thoroughly analyze the provided figure of the research paper, attaining a deep understanding of its contents, including but not limited to, the study’s objectives, methodologies, results, and conclusions.
-Develop 2 academic questions that: Must not mention the name of the figure directly or words like “figure”,“table” in the question.
-Demand complex reasoning and a comprehensive understanding of the whole figure.
-Must ask questions that can only be answered by the content in the figure, without reliance on textual information.
-Integrate knowledge from the entire figure, reflecting a nuanced understanding.
-Frame questions to elicit detailed and informed responses directly supported by the figure.
-Craft detailed answers for each question that: Are directly derived from the provided figure, excluding information not found within the material.
-Are comprehensive and cover all relevant aspects, as presented in the provided figure.
-Accurately reflect the information and insights offered by the research paper figure.
-Expected Output: List your Q and A as: [Q1]: [Formulate the first academic question here, ensuring it requires complex reasoning and encompasses the entirety of the provided figure for a comprehensive understanding.
-Based solely on the figure, without referencing accompanying text.
-Must not mention the name of the figure directly.
-Must not mention words like “figure”, “table” exc.] [A1]: [Provide a detailed answer here, derived exclusively from the information within the provided figure, ensuring the response is thorough and precise without including extraneous details.
-Mention from which figure/table you get the answer.] [Q2]: [A2]: Here is the supplementary paragraph text for the figure:
-Figure 7: Text-image QA generation prompt for the training set.
-Section QA Generation Prompt for the Training Set Task Definition: Generate 3 high-quality academic questions and corresponding answers based on a section of a research paper.
-Requirements: Each question should generate two types of answers.
-The first answer should be concise, directly addressing the question with minimal wording.
-The second answer should include a “chain of thought” that provides a reasoning process and be detailed.
-You should use English.
-Generate questions that: Can be answered with simple reasoning and only require a global understanding.
-The question should be able to answered directly with the material.
-Do not include 2 or more subquestions in each question.
-Ensure that the concise answers can be provided using sentences or phrases that do not exceed 20 words in length.
-Remember it is only a paragraph rather than a whole passage so do not ask global question about the study.
-Generate question-answer pairs instead of multiple-choice questions.
-Do not generate global question about the paper such as main idea or abstract.
-Must not mention the name of the figure directly or words like “figure”, “table” exc in the question.
-Must ask questions that can only be answered by the content in the figure, without reliance on textual information.
-Craft the 2 answers for each question that: Are directly derived from the provided figure, excluding information not found within the material.
-Are comprehensive and cover all relevant aspects, as presented in the provided figure.
-Accurately reflect the information and insights offered by the research paper figure.
-Expected Output: List your Q and A as: [Q1]: [Insert the academic question here, can be answered directly from the material] [A11]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words] [A12]: [Insert the detailed answer here, based solely on the given information.
-This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption.
-Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] [Q2]: [A21]: [A22]: [Q3]: [A31]: [A32]: Remember again all the questions should be academic and can be answered by the information in the material.Do not add extra information.Do not ask about the main idea.
-Here is the supplementary paragraph text for the figure:
-Figure 8: Section QA generation prompt for the training set.
-Cross-paragraph Question Generation Prompt for the Training Set Task Definition: Based on the selected paragraph from a research paper that share a thematic or conceptual connection, formulate an insightful, open-ended question.
-This question should reflect the shared themes or concepts of your selections and relate to the broader context of the research paper.
-Requirements: Ascertain the underlying connection among the paragraphs and the figures/tables(if provided).
-Subsequently, craft an insightful, open-ended question that encapsulates the identified themes or connections, aiming to foster analytical thinking and in-depth discussion on the subject matter of the paper.
-Note that your question should not directly include the “idx”s of the paragraphs.
-Expected Output: [Q]: [Your generated question based on the shared themes or information] Here are the selected paragraphs in the paper:
-Cross-paragraph Answer Generation Prompt for the Training set Task Definition: Given some selected paragraphs from a research paper, each chosen based on their relevance (excluding the first 3 paragraphs), and ensuring that these paragraphs share a certain level of association, you are to answer a question that is related to the content of these selected paragraphs.
-The question is crafted to encompass the themes or findings presented in the paragraphs of the chosen paragraphs, aiming for a comprehensive understanding and connection between these elements.
-Requirements: Craft the 2 answers for the question that: Are directly derived from the provided figure, excluding information not found within the material.
-Are comprehensive and cover all relevant aspects, as presented in the provided figure.
-Accurately reflect the information and insights offered by the research paper figure.
-Expected Output: [A1]: [Insert the concise answer here, providing a straightforward, brief response directly addressing the question, no more than 20 words.] [A2]: [Insert the detailed answer here, based solely on the given information.
-This answer must include a detailed “thought chain” or reasoning process, detailing how the conclusions are drawn from the image and caption.
-Must not mention the name of the figure/table from which the answer is derived directly, without adding extraneous details.] The question is:
-Figure 9: Cross-paragraph QA generation prompt for the training set.
-Text-only Question Generation Prompt for the Test set Task Definition: Create 2 academic questions from a given research paper paragraph.
-Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions.
-Focus on the paragraph,not the entire paper.
-If the paragraph lacks valid information,return ‘quit’.
-You should use English.
-Develop 2 questions that: No more than 30 words.
-Incorporate knowledge from the paragraph.
-Should be answered by text instead of one of the multiple choices.
-Elicit detailed responses supported by the text.
-Expected Output: (Return ‘quit’ directly if the paragraph lacks valid information.) [Q1]: question1 here [Q2]: question2 here
-Text-only Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: No more than one sentence, less than 20 words.
-Comprehensive and cover all relevant aspects.
-Accurately reflect the paragraph’s information and insights.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence.
-[thinking procedure]: [A2]: answer2 here, no more than one sentence.
-Text-only Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: Within a few keywords, less than 20 words.
-Comprehensive and cover all relevant aspects.
-Accurately reflect the paragraph’s information and insights.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords.
-[thinking procedure]: [A2]: answer2 here, within a few keywords.
-Figure 10: Text-only type QA generation prompt for the test set.
-Text-image Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on the provided figures and tables from a research paper.
-Requirements: Analyze the paragraph thoroughly,understanding its content including the study’s objectives, ethods, results,and conclusions.
-Focus on the paragraph,not the entire paper.
-If the paragraph lacks valid information,return ‘quit’.
-You should use English.
-Develop 2 questions that: No more than 30 words.
-Are specific to the unique data or details visible in the figures/tables and are answerable only based on the material without inferring or speculating on details not explicitly explained by the figures/tables.
-Must not mention the label of the figure/table directly or use words like ‘from the figure/table’.
-Expected Output: [Q1]: [Q2]:
-Text-image Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: No more than one sentence, less than 20 words.
-Always use English.
-Do not infer or speculate on details not explicitly explained by the figures/tables.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence.
-[thinking procedure]: [A2]: answer2 here, no more than one sentence.
-Text-image Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: Within a few keywords, less than 20 words.
-Comprehensive and cover all relevant aspects.
-Accurately reflect the paragraph’s information and insights.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords.
-[thinking procedure]: [A2]: answer2 here, within a few keywords.
-Figure 11: Text-image QA generation prompt for the test set.
-Section Question Generation Prompt for the Test set Task Definition: Formulate 2 academic questions based on a section from a research paper.
-Requirements: Carefully read and comprehend the entire provided section of the research paper to ensure a thorough understanding of its content, including key points, findings, methodologies, and conclusions.
-You should Always use English.
-Develop 2 questions that: No more than 30 words.
-Require an integration of information from all paragraphs and figures/tables in the section.
-Must not mention the label of the figure/table directly or use words like ‘from the figure/table’.
-Not based on common knowledge or assumptions not supported by the figures and tables.
-Expected Output: [Q1]: [Q2]:
-Section Answer Generation Prompt 1 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: No more than one sentence, less than 20 words.
-Always use English.
-Do not infer or speculate on details not explicitly explained by the figures/tables.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, no more than one sentence.
-[thinking procedure]: [A2]: answer2 here, no more than one sentence.
-Section Answer Generation Prompt 2 for the Test set Task Definition: Answer 2 questions based on the material given.
-Requirements: The answers should be: Within a few keywords, less than 20 words.
-Comprehensive and cover all relevant aspects.
-Accurately reflect the paragraph’s information and insights.
-You should think step by step and give you answer in the end of your generation like: [thinking procedure]: [A1/A2] Expected Output: [thinking procedure]: [A1]: answer1 here, within a few keywords.
-[thinking procedure]: [A2]: answer2 here, within a few keywords.
-Figure 12: Section QA generation prompt for the test set.
+Appendix D Limitation
+For our proposed dataset, current documents are mainly limited to academic papers, so the layout format and subject matter of the documents are relatively simple.
+We will expand the dataset with more diverse documents.
+Besides, our model is not specifically designed for some global queries, which will become a key research problem in the future.

### v3

@@
-{cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling
+Equal contribution Project lead
 Abstract
@@
 However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers.
-In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents.
+In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) that is designed to enhance multimodal question-answering (QA) for long PDF documents.
 PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM.
-The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model.
+The sparse sampler selects the paragraphs or diagrams most pertinent to user queries.
 To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers.
-Multiple strategies are proposed to automatically generate 1.1 million QA pairs along with their corresponding evidence sources.
+Multiple strategies are proposed to build high-quality 1.1 million QA pairs along with their corresponding evidence sources.
 Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1.
 Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
-footnotetext: Equal contribution; Corresponding author
 1 Introduction
-The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding [1, 2], where these models have demonstrated impressive capabilities in processing and generating human-like text.
+The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding saad2023pdftriage ; jacob2024constructing ; zhang2025smaller , where these models have demonstrated impressive capabilities in processing and generating human-like text.
 However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
@@
 1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables).
-They then employ long-context LLMs [3, 4, 5] or utilize Retrieval Augmented Generation (RAG) techniques [6, 7, 8, 9] to process the textual content.
+They then employ long-context LLMs peng2024yarn ; chen2024longlora ; NEURIPS2023_8511d06d or utilize Text Retrieval Augmented Generation (RAG) techniques yue2023disc ; tang2024multihop ; edge2024local to process the textual content.
 The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
 Figure 1: Method comparison for long multi-page PDF document understanding.
-(a) Plain text solution: long-context/RAG LLMs for parsed pure text content.
-(b) Pure vision solution: VDU models for page-level encoding and feature interaction.
-(c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
+(a) Plain text solution: long-context LLMs and TextRAG-based LLMs for parsed pure text content.
+(b) Pure vision solution: Multi-page MLLMs with page-level feature interaction and VisualRAG-based MLLMs for page images.
+(c) Our method is based on end-to-end multimodal sparse sampling for long PDFs with interleaved text and images.
+Note: “LLM” in (c) is the language backbone component integrated within MLLM.
 The second route is based on pure visual modality understanding.
-As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks[10, 11, 12].
+As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks liu2024textmonkey ; chen2024far ; wei2025vary .
 However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues.
-This makes it challenging to efficiently process multi-page long documents.
-While some recent multi-page VDU models handle up to 8 pages [13] or 20 pages [14], they typically encode each page separately and perform page-level visual feature interactions such as concatenation [15, 16].
-Nevertheless, as the number of pages grows, computational resource consumption still escalates substantially, rendering these models inefficient for processing longer documents.
+This makes it challenging to process multi-page long documents efficiently.
+While some recent multi-page visual document understanding models handle up to 8 pages liu2024focus or 20 pages TITO2023109834 , they typically encode each page separately and perform page-level visual feature interactions such as concatenation SlideVQA2023 ; blau2024gram .
+Nevertheless, as the number of pages grows, computational resource consumption escalates substantially, rendering these models inefficient for processing longer documents.
+Although some recent VisualRAG-based models yu2024visrag ; chen2025svrag try to find document pages related to the query, they perform poorly for text-intensive scenarios and cannot provide more fine-grained evidence.
 Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong.
@@
 It encodes text paragraphs and diagrams in the parsed PDF document, utilizing both text and image representations to identify and extract the most relevant evidence in response to a user’s query.
-The sampled sparse evidence significantly reduces the number of input tokens of LLM and this process is independent of the length of the input documents.
+The sampled sparse evidence significantly reduces the number of input tokens of the LLM and this process is independent of the length of the input documents.
 Moreover, the sparse sampler and LLM can be integrated in an end-to-end manner for training and inference, optimizing the performance of multimodal representation and question answering while improving time efficiency.
-It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
-Another important characteristic is that it can naturally provide strong interpretability for the question answering.
+It is worth noting that this sparse sampler is a flexible and generic design that can be applied to various MLLM architectures through fine-tuning.
+Another important characteristic is that it can naturally provide strong interpretability for question answering.
 In order to simultaneously represent and understand the multimodal content of documents and further improve the ability to process long PDF documents, we construct a training dataset specifically for English and Chinese academic paper PDFs.
@@
 We also provide a corresponding bilingual benchmark named PaperPDF.
-We train PDF-WuKong on PaperPDF dataset, complemented by general-domain document question-answering datasets.
+Table 1: Comparison of various models for processing multi-page long documents.
+Input modality Type Number of tokens Models
+Plain text Long-context Linear increase LongLoRA chen2024longlora , LongLLaMA NEURIPS2023_8511d06d , YaRN peng2024yarn
+Text RAG w/o Linear increase Graph RAG edge2024local , DISC-LawLLM yue2023disc , RAPTOR sarthi2024raptor
+Pure vision Multi-page Linear increase Fox liu2024focus , DocOwl2 hu2024mplugdocowl2 , MiniCPM-V yao2024minicpm ; yu2025minicpm , InternVL InternVL2.5 ; wang2025internvl3_5 , Qwen-VL Qwen2.5-VL ; Qwen3-VL , Granite Vision team2025granite , DocSeeker yan2026docseeker
+Visual RAG w/o Linear increase VisRAG yu2024visrag , SV-RAG chen2025svrag
+Text and images Multi-page Linear increase Hi-VT5 TITO2023109834 , GRAM blau2024gram
+Multimodal RAG w/o Linear increase PDF-WuKong (Ours)
+We train PDF-WuKong on the PaperPDF dataset, complemented by general-domain document question-answering datasets.
 Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding.
@@
 As the number of document pages increases, its accuracy and efficiency will not decrease significantly.
-It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17].
-Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters.
+It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE van2023document .
+Besides, for the recent benchmark MM-NIAH wang2024needle of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters.
 Our model achieves the best performance on multimodal content with a context length of 64K.
 The main contributions of this paper are as follows:
-  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient PDF question answering.
+  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient performance and providing explainable evidence for the answer.
   • We propose a bilingual PDF multimodal question answering dataset (PaperPDF) with $1.1M$ QA pairs for training and $10k$ QA pairs for evaluation.
   • Our model significantly surpasses existing open-source models and proprietary products (by an average of $8.6\%$ on F1) on long multimodal PDF understanding.
-Input modality Type Number of tokens Models
-Plain text Long-context Linear increase LongLoRA [4], LongLLaMA [5], YaRN [3]
-RAG w/o Linear increase Graph RAG [9], DISC-LawLLM [6], RAPTOR [19]
-Pure vision Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], XC2-4KHD [24], InternVL-V1.5 [11]
-Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [25], CREAM [26]
-Text and images Unlimited-page w/o Linear increase PDF-WuKong (Ours)
-Table 1: Comparison of various models for processing multi-page long documents.
 2 Related Works
 2.1 Document Understanding Datasets
-Early datasets focused on NLP tasks like summarization [27] and QA [28] of plain text, while visual document datasets targeted text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34].
-Recent multimodal document QA datasets include DocVQA [35] and OCRVQA [36] for single-page documents, ChartQA [37] and ChartX [38] focusing on visual reasoning in charts.
-Datasets like ArXivQA [39] and InfoVQA [40] enhance MLLMs’ abilities on academic and infographic documents.
-However, these datasets are limited to single-page tasks, and current MLLMs [41, 11] perform well on them.
-Multi-page QA datasets like MP-DocVQA [14], DUDE [17], DocGenome [42], and MM-NIAH [18] require understanding content relationships via multi-hop reasoning.
-Yet, answers in these datasets lack evidence and reliable interpretability, especially for questions needing multiple pieces of evidence from long documents.
+Earlier document understanding datasets only focused on the NLP tasks such as summarization huang2021efficient , and QA dasigi2021dataset of plain text.
+Meanwhile, there were several visual document datasets mainly aimed at text perception tasks such as Document Layout Analysis (DLA) zhong2019publaynet ; li2020docbank ; pfitzmann2022doclaynet and Key Information Extraction (KIE) simsa2023docile ; park2019cord ; huang2019icdar2019 .
+Recently, more datasets have been proposed for multimodal document QA across various scenarios.
+For example, DocVQA Mathew_2021_WACV and OCRVQA  MishraSSC19 provide single-page QA data on books and business documents.
+ChartQA masry2022chartqa and ChartX xia2024chartx focus on visual reasoning for chart documents.
+ArXivQA li2024multimodal extracts scientific figure-caption data from the arxiv papers to enhance the academic ability of MLLMs.
+InfoVQA Mathew_2022_WACV contains many infographic documents that are a combination of textual, graphical and visual elements.
+However, these visual document datasets only define the single-page VQA task.
+There are also some multi-page QA datasets SlideVQA2023 ; van2023document ; TITO2023109834 that require the model to understand the content relationship via multi-hop reasoning and capture the crucial information from multi-page documents.
+MP-DocVQA TITO2023109834 extends DocVQA Mathew_2021_WACV by adding the context pages.
+DUDE van2023document constructs multi-page QA data from multi-industry and multi-domain documents.
+Additionally, a large-scale dataset designed for the comprehensive understanding of lecture slides has been introduced zhang2025towards .
+DocGenome xia2024docgenome was constructed as a scientific document benchmark.
+MM-NIAH wang2024needle is a benchmark evaluating the capability of MLLMs to comprehend long multimodal documents, which requires the model to answer according to the key information scattered throughout the document.
+However, the answers in these datasets lack evidence and cannot provide reliable interpretability, especially for questions that require referring to multiple pieces of evidence in long documents.
 2.2 Document Understanding Methods
-Existing methods focus on plain text or limited document images.
-Text-based approaches convert documents into plain text using OCR and then employ long-context mechanisms like sparse attention [4], memory networks [5], or position interpolation [3].
-Retrieval-augmented generation methods [6, 19, 9] also handle long texts effectively.
-These approaches struggle with fine-grained visual understanding.
-Another solution, visual document understanding, treats each page as an image.
-MLLMs like UniDoc [20], mPLUG-DocOwl [21], and Vary [12] perform OCR-free understanding.
-Models such as UReader [22] and TextMonkey [10] divide high-resolution pages into patches.
-InternLM-XC2-4KHD [24] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration.
-However, reliance on high resolution increases token counts and isn’t scalable to multi-page documents.
-For multi-page documents, models like Hi-VT5 [14], GRAM [16], Fox [13], CREAM [26] and mPLUG-DocOwl2 [25] encode pages separately and perform page-level interactions.
-More pages generate more visual tokens, increasing resource consumption and inefficiency for longer documents.
-Thus, we propose parsing documents into interleaved text and images, followed by sparse sampling in an end-to-end manner.
+Existing document understanding methods typically focus on either plain text or a limited number of document images.
+Methods that rely on pure text modality aim to process documents by first converting them into plain text through document parsing or Optical Character Recognition (OCR) zhao2024cbnet ; zheng2024cdistnet ; wu2024end ; liu2025multi .
+They then employ efficient long-context mechanisms to handle long texts, such as sparse attention chen2024longlora , memory networks NEURIPS2023_8511d06d , or position interpolation peng2024yarn .
+Besides, the methods based on retrieval-augmented generation (RAG) yue2023disc ; sarthi2024raptor ; edge2024local also show impressive capabilities for long texts.
+While these approaches can integrate visual elements by including image captions or transforming images into natural language descriptions, they struggle with a fine-grained understanding of visual information.
+This restricts their effectiveness in tasks requiring detailed interpretation of textual and visual components within documents.
+Another solution is visual document understanding with purely visual input, treating each document page as an image.
+Many MLLMs such as UniDoc feng2023unidocuniversallargemultimodal , mPLUG-DocOwl ye2023mplugdocowl , StrucTexTv3 lyu2024structextv3 and Vary wei2025vary can perform this task in an OCR-free manner.
+Other researchers have advanced the understanding of high-resolution document pages by dividing input images into smaller patches, such as UReader ye2023ureader , TextMonkey liu2024textmonkey , and LLaVA-NeXT liu2024llavanext .
+DocPedia feng2023docpedia processes high-resolution images in the frequency domain via the DCT transformation.
+InternVL-V1.5 chen2024far introduces a dynamic resolution mechanism to capture more details.
+The reliance on high resolution results in a higher number of tokens and cannot be extended to multi-page documents.
+Recently, CoPali faysse2025colpali uses Vision LLM to retrieve query-related image patches in a single-page image, but it has not been applied to multi-page documents and only conducts the retrieval task without the QA task.
+There are several visual models specifically designed for multi-page documents.
+Fox liu2024focus unifies all image tokens of up to 8 pages into a sequence to achieve multi-page QA.
+mPLUG-DocOwl2 hu2024mplugdocowl2 compresses each high-resolution document image into 324 tokens, guided by low-resolution global visual features.
+Notably, DocSeeker yan2026docseeker embeds the retrieval capability directly into the model rather than relying on a separate RAG pipeline, introducing structured visual reasoning to achieve accurate and fine-grained evidence grounding in multi-page documents.
+To improve model interpretability, DocThinker yu2025docthinker leverages rule-based reinforcement learning for explainable document understanding.
+Besides, some general MLLMs such as MiniCPM-V yao2024minicpm ; yu2025minicpm , InternVL InternVL2.5 ; wang2025internvl3_5 , Qwen-VL Qwen2.5-VL ; Qwen3-VL , and Granite Vision team2025granite can already support multi-image input by training with multi-image samples.
+These models encode each page image separately and then perform page-level visual feature interactions such as concatenation.
+However, more pages generate more visual tokens, which brings greater resource consumption to the LLM.
+This makes them inefficient and unable to process longer documents.
+Recently, some document understanding MLLMs based on pure visual RAG yu2024visrag ; chen2025svrag ; yinmstar have been proposed.
+They usually separate the page image retrieval and document understanding generation into two independent stages during training and inference.
+They still treat each page as a separate image and use pure visual retrieval to obtain page-level evidence.
+These retrieval schemes remain challenging for text-intensive document understanding due to insufficient text recognition capabilities.
+Moreover, they can only obtain page-level evidence rather than more fine-grained evidence.
+Beyond plain text and pure vision approaches, certain specialized models for multi-page documents attempt to integrate multimodal information by concatenating OCR tokens with visual features for each page.
+Hi-VT5 TITO2023109834 and GRAM blau2024gram are representative works in this category.
+They combine tokens of multiple pages through hierarchical transformer architecture TITO2023109834 or global-local reasoning blau2024gram , being able to handle up to 20 pages.
+However, their token count still increases linearly with the number of pages.
+Therefore, we propose first to parse the lengthy document into interleaved text and image content and then perform multimodal sparse sampling in an end-to-end manner.
 Tab.
-1 summarizes these methods.
+1 summarizes the attributes of each type of method.
 3 Methodology
-Figure 2: The overall structure of PDF-WuKong consists of a document parser, a sparse sampler and a large language model.
+Figure 2: The overall structure of PDF-WuKong consists of a document parser, a multimodal sparse sampler that integrates end-to-end with an LLM.
+“LLM” in the figure refers to the language backbone within the MLLM.
+3.1 Motivation and Method Overview
+The central motivation behind PDF-WuKong is to empower multimodal large language models (MLLMs) to efficiently deal with sophisticated multimodal reasoning and comprehension tasks on lengthy multimodal PDF documents, particularly academic papers densely populated with paragraphs, figures, and tables.
+Directly inputting entire documents, including all textual paragraphs, figures, and tables, into existing MLLMs is computationally prohibitive due to their substantial length.
+Furthermore, crucial information relevant to a user’s query may be diluted by irrelevant content, making it challenging for the model to accurately locate and reason about key evidence.
+To address these challenges, our overarching insight is to leverage sparse multimodal sampling to selectively identify and retrieve the most relevant text and image blocks given a user query.
+Crucially, instead of introducing a fully independent multimodal retrieval model, we propose a lightweight retrieval strategy that reuses the visual encoder from the pretrained MLLM.
+This strategy reduces computational redundancy and facilitates efficient embedding caching, enabling fast inference and effective end-to-end training.
+As illustrated in Fig.
+2, PDF-WuKong comprises three main components: 1) Document Parsing converts PDF documents into structured, interleaved sequences of text and image blocks (Sec.
+3.2); 2) Multimodal Sparse Sampling encodes and caches compact embeddings for retrieval, efficiently identifying query-relevant sparse subsets of blocks (Sec.
+3.3); 3) Multimodal Response Generation takes the retrieved sparse set of multimodal blocks and synthesizes concise, accurate responses to user queries using a pretrained MLLM (Sec.
+3.4).
+The overall inference process is summarized in Algorithm 1.
+3.2 Document Parsing and Representation
+Given a PDF document $D$, we first parse it into a structured sequence of multimodal blocks, explicitly preserving its layout and reading order. Specifically, we extract:
+- A set of text blocks $\{T_{1},T_{2},\dots,T_{n}\}$, each typically corresponding to a paragraph.
+- A set of image blocks $\{I_{1},I_{2},\dots,I_{m}\}$, each including figures or tables with their captions.
+If an image block has an accompanying textual caption, we explicitly associate them together.
+The structured multimodal content is stored as an XML file and parsed offline using open-source PDF parsing tools Grobid GROBID and MinerU wang2024mineruopensourcesolutionprecise .
+To ensure robust parsing of complex academic layouts, we employ a hybrid parsing pipeline combining Grobid GROBID for metadata extraction and MinerU wang2024mineruopensourcesolutionprecise for fine-grained layout analysis.
+We selected this specific combination after comparing multiple open-source alternatives.
+First, both tools are lightweight and computationally efficient, enabling the scalable processing of our massive 1.1M document corpus.
+Second, MinerU is specifically optimized for academic documents; it demonstrates superior performance in parsing complex elements such as mathematical formulas and charts compared to traditional rule-based parsers.
+Crucially, its robust support for Chinese character recognition ensures high-quality parsing for the Chinese subset of our bilingual dataset.
+This pipeline effectively mitigates common parsing errors in mixed-content documents.
+3.3 Multimodal Sparse Sampling
+To effectively handle lengthy documents, our sparse sampling approach aims to identify a concise subset of the most query-relevant text and image blocks.
+To achieve this, we encode each block into a compact embedding vector, enabling efficient retrieval through similarity matching in a unified embedding space.
+Text Embedding.
+We introduce a lightweight text encoder $E_{T}(\cdot)$ for each text block $T_{i}$, resulting in a single text embedding vector $e_{T_{i}}\in\mathbb{R}^{d_{T}}$ that concisely captures semantic content.
+Image Embedding.
+For each image block $I_{j}$, we reuse the visual encoder $E_{V}(\cdot)$ from the pretrained MLLM to obtain the visual sequence embeddings $e_{I_{j}}\in\mathbb{R}^{n\times d_{V}}$, where $n$ is the number of visual tokens and $d_{v}$ is the visual embedding dimension.
+To unify visual and textual embeddings into a shared embedding space, we aggregate the visual sequence into a single compact vector and project it into the textual embedding space $d_{t}$ using the projection head $MLP_{proj}$}:
+$e_{I_{j}}^{\prime}=MLP_{proj}(MeanPool(e_{I_{j}})),\quad e_{I_{j}}^{\prime}\in\mathbb{R}^{d_{T}}.$} (1)
+If an image block has an associated textual caption, we further enrich its embedding by encoding the caption text using the same text encoder $E_{T}(\cdot)$ and obtain a single caption embedding vector.
+The final embedding $e_{I_{j}}^{\text{fused}}$ for the image is obtained by element-wise addition of the projected visual embedding and the caption embedding.
+Query-driven Retrieval.
+Given a user query $q$, we also encode it using the text encoder $E_{T}(\cdot)$ to obtain a compact query embedding vector in $e_{q}\in\mathbb{R}^{d_{T}}$.
+Then, we can compute cosine similarity between the query embedding and all cached compact embeddings of text blocks ($\{e_{T_{i}}\}$) and image blocks ($\{e_{I_{j}}^{\text{fused}}\}$).
+The top-$K$ most similar blocks are selected as sparse evidence for multimodal response generation.
+Offline Caching.
+All compact embeddings for text and image blocks, along with the original visual sequence embeddings, are computed offline and cached.
+This offline caching strategy significantly speeds up inference by avoiding redundant visual encoding computations during response generation.
+3.4 Multimodal Response Generation
+Given the retrieved sparse multimodal blocks and the user query, we generate the final response through the pretrained MLLM.
+For image blocks, we reuse the cached visual sequence embeddings $\{e_{I_{j}}\}$, containing all image patch tokens.
+Since these are derived from the MLLM’s native visual encoder, they can be directly projected into the input space of the MLLM’s language backbone using the original connection module $MLP_{conn}$, preserving fine-grained visual details:
+$\hat{e_{I_{j}}}=MLP_{conn}(e_{I_{j}}),\quad\hat{e_{I_{j}}}\in\mathbb{R}^{n\times d_{llm}},$} (2)
+$\hat{e_{I_{j}}}$ serves as the visual context tokens fed into the LLM backbone. This approach eliminates redundant visual encoding operations at inference time.
+Algorithm 1 Inference pipeline for PDF-WuKong
+1:Input: PDF document $D$, user query $q$
+2:Output: Generated answer $a$
+3:Initialize: Text encoder $E_{T}$, image encoder $E_{V}$, LLM, and two MLP projectors $\text{MLP}_{\text{proj}}$, $\text{MLP}_{\text{conn}}$
+4:Stage 1: Document Parsing
+5:Parse the input document $D$ into text blocks and images with their captions:
+$\{T_{i}\},\{(I_{j},C_{j})\}\leftarrow\text{Parser}(D)$
+6:Stage 2: Sparse Sampling
+7:Encode all text blocks and images and cache all candidate vector embeddings:
+$\{e_{T_{i}}\}\leftarrow E_{T}(\{T_{i}\}),$
+$\{e_{I_{j}}\}\leftarrow E_{V}(\{I_{j}\}),$
+$\{e_{C_{j}}\}\leftarrow E_{T}(\{C_{j}\}),$
+$\{e_{I_{j}}^{\prime}\}\leftarrow\text{MLP}_{\text{proj}}(MeanPool(\{e_{I_{j}}\})),$}
+$\{e_{I_{j}}^{\text{fused}}\}\leftarrow\{e_{I_{j}}^{\prime}\}+\{e_{C_{j}}\}$
+8:Encode the user query $q$:
+$e_{q}\leftarrow E_{T}(q)$
+9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{e_{T_{i}}\}$, $\{e_{I_{j}}^{\text{fused}}\}$:
+$S_{T}=\{\text{sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
+$S_{I}=\{\text{sim}(e_{q},e_{I_{j}}^{\text{fused}})\mid j=1,2,\dots,m\}$
+10:Select top-$k$ relevant text blocks and images with captions:
+$(T,I,C)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$}
+11:Stage 3: Answer Generation
+12:Input the query $q$ and the selected top-$k$ text blocks $T_{i}$ and visual embeddings $e_{I_{j}}$ into the LLM:
+$\{\hat{e_{I_{j}}}\}\leftarrow\text{MLP}_{\text{conn}}(\{e_{I_{j}}\}),$
+$a\leftarrow\text{LLM}(q,(T_{i},\hat{e_{I_{j}}},C_{i})_{top})$}
+13:Return the generated answer $a$.
+Conversely, for text blocks, we provide raw textual input.
+Reusing retrieval embeddings is not feasible because 1) the text encoder (BGE-M3) and the MLLM’s language module use different tokenizers and vocabularies, and 2) LLMs are optimized to process discrete token indices mapped to their own internal embedding spaces.
+Providing raw text ensures the model utilizes its pre-trained linguistic capabilities to the fullest extent.
+It is worth noting that we utilize the original visual embeddings $\{e_{I_{j}}\}$ for generation rather than the fused embeddings $\{e_{I_{j}}^{fused}\}$ used in retrieval.
+This decision is based on the fact that the $MLP_{conn}$ is pre-trained to align pure visual features with the LLM space; introducing fused embeddings at this stage would cause a distribution mismatch.
+To encourage transparent reasoning and enhance semantic understanding, we explicitly annotate each input block with metadata such as block indices and section headings.
+Specifically, for retrieved image blocks, their corresponding captions are also provided as plain text inputs to the LLM alongside the visual embeddings.
+The MLLM is trained to explicitly reference relevant block indices when generating responses, thereby enhancing interpretability and evidence grounding.
+3.5 Joint End-to-End Training Objectives
+PDF-WuKong is trained end-to-end using a combination of two complementary learning objectives: (1) a contrastive representation learning loss $\mathcal{L}_{\text{CTR}}$ for effective embedding alignment, and (2) a cross-entropy loss $\mathcal{L}_{\text{CE}}$ for accurate multimodal response generation.
+Contrastive Representation Loss ($\mathcal{L}_{\text{CTR}}$).
+For each training instance (query-answer pair), positive blocks ($B^{+}$) are those explicitly annotated as relevant evidence, while negative blocks ($B^{-}$) are randomly sampled irrelevant blocks from the document.
+Given a query embedding $e_{q}\in$, we optimize embeddings by minimizing the following contrastive loss:
+$\mathcal{L}_{\text{CTR}}=-\frac{1}{|B^{+}|}\sum_{e^{+}\in B^{+}}\log\frac{\exp\bigl(\text{sim}(e_{q},e^{+})/\tau\bigr)}{\mathcal{Z}},$ (3)
+where the normalization term $\mathcal{Z}$ sums over positive and negative samples:
+$\mathcal{Z}=\exp\bigl(\text{sim}(e_{q},e^{+})/\tau\bigr)+\sum_{e^{-}\in B^{-}}\exp\bigl(\text{sim}(e_{q},e^{-})/\tau\bigr).$
+Here, $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a temperature hyperparameter controlling the softness of the similarity distribution.
+Cross-Entropy Response Generation Loss ($\mathcal{L}_{\text{CE}}$).
+The second objective trains the MLLM to generate accurate textual responses conditioned on the retrieved multimodal evidence and the query:
+$\mathcal{L}_{\text{CE}}=-\sum_{l=1}^{L}\log p_{\theta}(r_{l}^{*}|q,B^{+},r_{<l}^{*}),$ (4)
+where $r^{*}$ is the ground-truth response sequence, and $p_{\theta}$ is the MLLM.
+Joint Optimization. Our end-to-end training objective integrates both losses:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CTR}}+\mathcal{L}_{\text{CE}}.$ (5)
+Overall, PDF-WuKong introduces joint alignment to reconcile efficiency and fidelity: the visual features first map to a text-dominated metric space for coarse cross-modal retrieval, then to the LLM-aligned compositional space for fine-grained reasoning.
+This design mirrors human coarse-to-fine comprehension, filtering salient evidence via unified semantics before synthesizing grounded responses.
+Meanwhile, it avoids redundant recomputation through parameter-efficient adaptation and offline caching.
+The duality resolves modality granularity conflicts, ensuring computational tractability without sacrificing pretrained multimodal synergies.
+4 PaperPDF Dataset
 4.1 Overview
-Our pipeline consists of three components: a document parser, a sparse sampler, and a large language model, as shown in Fig.
-2.
-The document parsing stage converts input PDFs into machine-readable content with interleaved text and images.
-The sparse sampler then encodes and caches embeddings for text blocks and images separately.
-When receiving a user query, it retrieves the most relevant content through similarity matching.
-Finally, the query and sampled tokens are fed into the LLM for answer generation.
-The detailed procedure is outlined in Algorithm 1 in the appendix.
-3.2 Document Parsing
-Given a PDF document $D$, the goal of document parsing is to convert it into some machine-readable text blocks $\{T_{1},T_{2},\dots,T_{n}\}$ and diagrams $\{I_{1},I_{2},\dots,I_{m}\}$ according to the reading order and layout structure.
-By default, text blocks are organized into paragraphs, and all figures and tables are saved as images.
-These text and images are finally reorganized into an XML file in reading order.
-This process can be completed using existing open-source PDF parsing tools.
-During inference, we directly input the parsed structured full data into the subsequent stage of PDF-WuKong.
-3.3 Sparse Sampling
-For a lengthy multi-page document, if it is directly input into the LLM, there will be two problems.
-The first is the problem of computing efficiency.
-The consumption of computing resources will increase dramatically.
-The second is the problem of inaccurate attention.
-Key information related to the user query is easily submerged by a large amount of irrelevant content.
-It is difficult for the model to accurately locate and extract important information in a huge token sequence.
-Therefore, sparse sampling is essential for efficiently handling lengthy multi-page documents by identifying and extracting the most relevant text chunks or diagrams based on their similarity to the user query.
-During the training, for the parsed $n$ text chunks $\{T_{1},T_{2},\ldots,T_{n}\}$, $m$ images $\{I_{1},I_{2},\ldots,I_{m}\}$, and the input user query $q$, we first extract the positive samples and the negative samples for the query.
-Our PaperPDF dataset has provided corresponding positive single-evidence or multi-evidence samples for each query-answer pair (detailed in Sec.
-4).
-We randomly select two text blocks and two images from the remaining text blocks and images as negative samples.
-Then, we use a text encoder $En\_T$ to obtain the text embeddings $e_{T_{P}},e_{T_{N}}$ and the query embedding $e_{q}$.
-An image encoder $En\_I$ is utilized to output the image features $e_{I_{P}},e_{I_{N}}$, which is shared with MLLM.
-Given the embeddings of the user query $e_{q}$, the positive samples $E_{P}=\{e_{T_{P}},e_{I_{P}}\}$ and negative samples $E_{N}=\{e_{T_{N}},e_{I_{N}}\}$, we employ a contrastive learning approach to align the text and image features with the query.
-The goal is to enable the model to capture the document content that is most relevant to the query.
-The contrastive learning loss is:
-$\mathcal{L}_{\text{rep}}=-\frac{1}{P}\sum_{e_{i}\in E_{P}}\log\frac{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}}{e^{\frac{\text{sim}(e_{q},e_{i})}{\tau}}+\displaystyle\sum_{e_{j}\in E_{N}}e^{\frac{\text{sim}(e_{q},e_{j})}{\tau}}},$ (1)
-where $\text{sim}(e_{q},e_{i})$ and $\text{sim}(e_{q},e_{j})$ represent the similarity between the query and the positive/negative samples.
-$\tau$ is the temperature parameter that controls the scale of the similarity scores.
-$P$ is the number of positive samples.
-By maximizing this probability, the model encourages the representations of the query and positive samples to be closer while pushing the representations of the query and negative samples apart.
-It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs.
-During the inference, we pre-encode all text blocks and images and cache all candidate embeddings.
-When the user inputs a query, we calculate the similarity between query embedding and cached text/image embeddings.
-Then the model automatically selects the top-$k$ relevant text blocks and images as evidence to respond to this query.
-Therefore, this process samples out sparse document content, greatly reducing the computational burden of the subsequent LLM and alleviating the problem of attention shift when facing ultra-long sequences.
-Moreover, the multimodal embedding cache further optimizes inference time.
-The pseudocodes for training and inference are shown in the appendix.
-3.4 Answer Generation
-At this stage, the large language model only receives the document content that is most relevant to the query and discards a lot of redundant information, so it can generate more accurate answers with higher efficiency.
-Specifically, we input the sampled top-$k$ evidence, the user query, and the task instruction into the LLM, and let it generate an answer based on the provided query and evidence.
-Considering that MLLM needs to encode images first for multimodal understanding, we directly input the image tokens obtained from the sparse sampler into the LLM, to save one image encoding process.
-Thus, the sparse sampler shares the same vision encoder with the MLLM.
-They can be integrated and trained in an end-to-end manner.
-During the training, we input the positive text $T_{P}$ and the positive image tokens $e_{I_{P}}$ into the LLM.
-Besides, the query and instruction are also input into the LLM.
-Then, we calculate the cross-entropy loss $\mathcal{L}_{\text{QA}}$ between the output answer $a$ and the ground truth.
-Finally, the total optimization objective is:
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}.$ (2)
-PDF-WuKong is optimized end-to-end by these two loss functions for effective multimodal alignment and QA.
-4 PaperPDF Dataset
+In order to enable PDF-WuKong to simultaneously represent and understand the multimodal content of documents and provide explainable answer evidence, we construct a dataset with English and Chinese academic paper PDFs.
+The dataset contains complete PDF documents, professional academic questions, answers, and evidence sources for the answers, based on multiple construction strategies.
+Specifically, our motivation for creating PaperPDF is threefold: (1) In the long document QA task, answers often derive from specific document segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multi-page multimodal document QA dataset.
+Therefore, we introduce multiple strategies to construct question-answer pairs from long documents for training and evaluation.
 Figure 3: The construction process of PaperPDF based on single evidence and multiple evidence.
-4.1 Overview
-Our motivation for creating PaperPDF is threefold: (1) In long document QA contexts, answers often derive from specific segments, with other content acting as noise and complicating MLLM reasoning; (2) Existing datasets are either limited to single-page documents or lack fine-grained evidence ground truth, hampering the training of our sparse sampler; (3) There is currently a lack of a bilingual multimodal PDF QA dataset.
-Therefore, we introduce a method for automatically generating question-answer pairs from long documents and present PaperPDF, a dataset designed for both training and evaluation.
-4.2 Dataset construction strategy
-The PaperPDF dataset is constructed through a four-step process as shown in Fig.
-3: document parsing, evidence extraction, QA generation, and data filtering.
-First, $100k$ PDFs are parsed to extract text and image chunks.
-Then we extract some evidence from each PDF according to our specific rules.
-Given the evidence and the generation prompt as the input, we use Gemini Pro [43] to generate questions and answers for the training set due to its free and rapid accessibility.
-GPT-4V [44] is used for the test set to ensure high evaluation quality.
-After obtaining the preliminary dataset, we conduct automatic filtering and remove abnormal data.
-For the test data, we further perform manual filtering from multiple aspects to ensure its quality.
-According to different evidence extraction rules, the triplets in PaperPDF can be divided into two categories.
-Single-evidence Q-E-A triplets.
-The questions in these triplets can be answered based on a single text chunk or a diagram.
-Therefore, the evidence can be categorized into Text-only and Image-only.
-These triplets enable PDF-WuKong to initially acquire the capabilities of sparse sampling and long multimodal document understanding.
-Multi-evidence Q-E-A triplets.
-These triplets require reasoning across multiple chunks, involving combinations of text and images.
-The types include Image-text (derived from a text chunk and associated images), Section (generated from all chunks in a section), and Cross-paragraph (involving related paragraphs across a document).
-For the third type, semantic summarizations for each paragraph are conducted first, followed by a selection of several related text chunks for QA generation.
-These multi-evidence triplets enhance our model’s multi-hop reasoning capability.
-In total, we obtained $1.1M$ bilingual training data and $10k$ testing data. The dataset statistics are shown in Tab. 2. The appendix contains more data statistics.
+4.2 Dataset construction
+The academic paper PDF is a kind of typical document that contains rich interleaved text and images, which can intuitively reflect the challenges of our task and the advantages of our model.
+Therefore, we first obtained approximately 60k English and 10k Chinese documents from open-access academic paper websites, encompassing nearly 70 disciplines, such as computer science, engineering, and materials science.
+To build high-quality data, we propose a four-step construction strategy and design a variety of evidence distributions based on human questioning habits.
+The construction strategy consists of document parsing, evidence extraction, QA generation, and data filtering, as shown in Fig.
+3.
+First, each PDF is parsed as interleaved text and image chunks according to the statment in Sec. 3.2.
+Second, we extract evidence from each PDF according to different extraction rules, and the type of evidence in our PaperPDF can be divided into two categories:
+Single-evidence extraction.
+The evidence can be categorized into Text-only and Image-only.
+We randomly extract a text paragraph or a diagram from a document and construct questions and answers based only on the information in this single evidence.
+We stipulate that the length of the extracted paragraphs must contain at least 3 sentences, and the diagram needs to be extracted together with its caption.
+These enable PDF-WuKong to acquire the initial capabilities of sparse sampling and long multimodal document understanding.
+Multi-evidence extraction.
+The evidence can be categorized into Image-text, Section, and Cross-paragraph.
+These facilitate reasoning across multiple chunks, involving combinations of text and images.
+The Image-text evidence is derived from a text chunk and the diagrams it refers to.
+It is also extracted from a diagram and the text paragraphs that reference it.
+The Section evidence utilizes all chunks in a section, including all the text paragraphs, figures, and tables.
+The Cross-paragraph evidence involves related paragraphs across the document.
+To extract this type of evidence, we first conduct semantic summarizations for each paragraph and then select several of the most semantically relevant text chunks.
+These multi-evidence samples can significantly enhance our model’s multi-hop reasoning capability.
+Third, we leverage commercial MLLM APIs reid2024gemini ; 2023gpt4V to generate questions and answers by feeding the API with extracted evidence and carefully engineered prompts.
+We implement specific prompts for different types of evidence, including the task definition, requirements and expected output.
+For question generation, the task definition is to let it generate two academic questions based on the given evidence and require the generation process to rely only on the provided content and eliminate other knowledge.
+Questions must be brief and objective and fully account for the provided evidence.
+For answer generation, the model requires thinking step by step based on the question and evidence and generating two answers.
+The expected outputs include a long answer and a short answer, each with its own thought process.
+Short answers are used for training and evaluation of PDF-WuKong.
+Long answers and thought processes can be used to further optimize models to have document reasoning capabilities in the future.
+For detailed prompt engineering and the Query-Evidence-Answer (Q-E-A) triplet examples of each type, please refer to the Appendix A.
+Given the evidence and the prompt as the input, we use Gemini Pro reid2024gemini to generate questions and answers for the training set due to its free and rapid accessibility.
+GPT-4V 2023gpt4V is used for the test set to ensure high evaluation quality.
+Using different APIs causes data bias but poses a higher challenge to the generalizability of the model.
+Table 2: The statistics of QA pairs for various types of evidence in English and Chinese.
 Category Train (En/Zh) Test (En/Zh)
@@
 Cross-paragraph 1.2k/0 118/0
-Table 2: The statistics of PaperPDF in English and Chinese.
+Table 3: Statistical analysis compared with other multi-page document datasets. Page and Token represent the average number of pages per document, and the average number of OCR tokens per document.
+MPdataset Doc. QA Lang. Page Token
+DUDE van2023document 5k 41.5k En 6 1831
+MP-Docvqa TITO2023109834 6k 46k En 8 2026
+SlideVQA SlideVQA2023 2.6k 14.5k En 20 1489
+PaperPDF (En) 60k 900k En 25 11371
+PaperPDF (Zh) 10k 200k Zh 11 3413
+PaperPDF 70k 1.1M En+Zh 23 10234
+Fourth, to ensure data quality, we conduct data filtering for the Q-E-A triplets.
+We designed many effective rules to automatically eliminate low-quality samples and remove abnormal data.
+For example, too-short questions or too-long answers may confuse the model, and duplicate questions will reduce diversity.
+After performing automatic filtering for the training and testing data, we performed manual filtering and correction for the testing data to improve the quality further.
+This human verification step is crucial to mitigate potential biases inherent in model-generated content.
+By having human experts verify and correct the Q-E-A triplets, we ensure that the test set ground truth reflects accurate human reasoning rather than the stylistic patterns of a specific commercial API.
+We checked if the questions made sense, if the answers were correct, if the Q-E-A triplets were related to each other, etc.
+We will present the statistical analysis and quality verification of the dataset in the next section.
+4.3 Dataset analysis
+In total, we obtained $1.1M$ bilingual training data and $10k$ testing data.
+The number of QA pairs for various types of evidence is shown in Tab.
+2.
+To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis in Tab.
+3.
+Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents and QA pairs but also features more pages and OCR tokens.
+More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images.
+Figure 4: Statistics of question types in PaperPDF dataset.
+Fig.
+4 shows the distribution of question types in our benchmark.
+We basically divide the types of questions in multimodal PDF documents into 7 categories, covering various questions that users may encounter.
+In order to further verify the correctness of the answers corresponding to each type of question, we manually evaluated the answers for each type in the test set.
+We first randomly selected 100 samples from each type and then asked two doctoral students with scientific research training to give an answer to each question.
+If the two people’s answers disagree, they need to discuss with each other to determine the final answer.
+Finally, we counted the consistency between the answers from the test set and the answers given by humans, as shown in Tab.
+4.
+The high agreement in the answers demonstrates the high quality of our benchmark.
+Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
+Explicit constructing Q-E-A triplets aligns with PDF-WuKong’s dual training objectives and ensures PDF-WuKong learns to distinguish fine-grained evidence patterns essential for long-document QA.
+Table 4: Answers’ consistency from the testing dataset and human evaluation for different types of questions.
+Question Type Consistency
+Single-hop understanding 95.2%
+Single-hop logical reasoning 93.0%
+Multi-hop understanding 93.6%
+Multi-hop logical reasoning 92.3%
+Figure understanding 94.8%
+Table understanding 97.3%
+Formula understanding 96.1%
 5 Experiments
+Table 5: Performance comparison with open-source models for long PDF understanding on PaperPDF.
+* indicates that the RAG model has been trained on our dataset.
+2p and 3p indicate whether VisRAG retrieves 2 or 3 pages of evidence.
+The best results are marked bold and the second results are underlined.
+Model #Param English Chinese Token
+ANLS F1 Rouge GPT-Acc ANLS F1 Rouge GPT-Acc
+Plain Text Solution
+IXC2-VL internlmxcomposer2 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
+IXC2-VL internlmxcomposer2 +BGE-M3 chen2024bge 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
+InternVL2 chen2024far 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
+InternVL2 chen2024far +BGE-M3 chen2024bge 8.5B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
+InternVL2 chen2024far +BGE-M3* chen2024bge 8.5B 36.4 37.3 34.7 53.3 28.4 34.2 34.5 46.3 849
+Pure Vision Solution
+Hi-VT5 TITO2023109834 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
+IXC2-VL internlmxcomposer2 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
+InternVL2 chen2024far 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
+Qwen2.5-VL Qwen2.5-VL 7B 35.5 34.8 32.6 38.3 23.5 34.4 38.7 40.8 5878
+MiniCPM-V 2.6 yao2024minicpm + VisRAG* yu2024visrag (2p) 10B 31.8 31.0 27.2 27.6 20.5 27.0 30.1 35.9 1444
+MiniCPM-V 2.6 yao2024minicpm + VisRAG* yu2024visrag (3p) 10B 32.4 32.5 28.5 28.4 20.8 27.9 30.8 37.7 2199
+Parsed Interleaved Text and Images
+IXC2-VL internlmxcomposer2 8B 27.8 31.2 32.6 37.7 22.5 29.5 29.2 31.4 6217
+InternVL2 chen2024far 8B 33.4 36.2 36.6 54.3 28.5 40.6 42.0 54.7 6220
+PDF-WuKong (ours) 8.5B 41.9 43.5 40.9 77.5 40.9 47.8 48.6 57.8 2107
 5.1 Implementation Details
-The LLM and the vision encoder are initialized from IXC2-VL-4KHD [24] and the maximum number of dynamic tiles is set as 16.
-The text encoder is initialized from BGE-M3 [45].
-We train the model by leveraging several document datasets including PaperPDF, DocVQA [35], ChartQA [37], InfoVQA [40], MPDocVQA [46], and DUDE [17].
-Before both training and testing, PaperPDF is parsed using Grobid [47] and MinerU [48], while the other datasets are processed following their default instructions.
+The LLM and the vision encoder are initialized from IXC2-VL-4KHD internlmxcomposer2_4khd and the maximum number of dynamic tiles is set as 16.
+The text encoder is initialized from BGE-M3 chen2024bge .
+We train the model by leveraging several document datasets including PaperPDF, DocVQA Mathew_2021_WACV , ChartQA masry2022chartqa , InfoVQA Mathew_2022_WACV , MPDocVQA TITO2023109834 , and DUDE van2023document .
+Before both training and testing, PaperPDF is parsed using Grobid GROBID and MinerU wang2024mineruopensourcesolutionprecise to support sparse sampling.
+In contrast, for the other datasets, we utilize the complete full-page images paired with their corresponding OCR text as global input blocks, bypassing the fine-grained parsing step.
 The training is conducted for one epoch using 128 Ascend 910B NPUs with a learning rate of 4e-5.
@@
 They contain 488 and 317 QA samples, respectively.
-Model #Param English Chinese Token
-ANLS F1 Rouge GPT-Acc ANLS F1 Rouge GPT-Acc
-Plain Text Solution
-IXC2-VL [49] 8B 27.4 30.8 32.8 37.6 21.1 28.5 28.7 30.8 4644
-IXC2-VL-RAG [49] 8.5B 32.4 34.0 32.4 48.4 23.3 29.8 32.3 38.1 623
-InternVL2 [11] 8B 19.5 28.0 27.6 37.5 24.4 28.2 27.7 30.8 4051
-InternVL2-RAG [11] 8B 29.8 29.8 28.3 50.2 24.6 27.8 26.8 43.0 583
-Pure Vision Solution
-Hi-VT5 [46] 0.3B 13.5 3.1 3.7 15.2 - - - - 11589
-IXC2-VL [49] 8B 27.1 24.8 25.1 20.5 15.9 19.5 22.2 27.4 4712
-InternVL2 [11] 8B 29.4 33.1 35.0 35.5 20.3 33.9 37.8 37.7 5008
-Parsed Interleaved Text and Images
-IXC2-VL [49] 8B 27.8 31.2 32.6 37.7 22.5 29.5 29.2 31.4 6217
-InternVL2 [11] 8B 33.4 36.2 36.6 54.3 28.5 40.6 42.0 54.7 6220
-PDF-WuKong (ours) 8.5B 41.9 43.5 40.9 77.5 40.9 47.8 48.6 57.8 2107
-Table 3: Performance comparison with open-source models for long PDF understanding on PaperPDF. The best results are marked bold and the second results are underlined.
-Model ANLS F1 Rouge GPT-Acc
-Gemini pro [50] 26.6 29.0 29.8 67.9
-Kimi [51] 28.5 33.6 31.1 74.7
-ChatGLM [52] 31.2 35.4 32.0 73.5
-Qwen [53] 36.0 40.3 35.5 78.1
-PDF-WuKong (ours) 41.8 43.2 40.7 77.5
-Table 4: Performance comparison with commercial products (tested in Sep 2024) for long PDF understanding.
-The results are tested on a subset of 50 English PDFs.
-Note: These are the products based on the models rather than the models themselves.
+Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products.
+The red box indicates the evidence that the correct answer depends on.
+Our model can find the text chunks, figures or tables contaning the evidence and output correct answers.
 5.2 Long PDF Understanding
 To assess the effectiveness of our model in understanding long PDF documents, we conduct comprehensive experiments comparing it with both open-source models and commercial products on the PaperPDF dataset.
-Due to limited capabilities of traditional document understanding models  [46, 16, 54, 55, 56] and the huge resource costs in advanced MLLMs  [49, 24, 11], achieving deep understanding of lengthy PDF documents remains a highly challenging task.
+Due to the limited capabilities of traditional document understanding models  TITO2023109834 ; blau2024gram ; huang2022layoutlmv3 ; beltagy2020longformer ; zaheer2020big and the huge resource costs in advanced MLLMs  internlmxcomposer2 ; internlmxcomposer2_4khd ; chen2024far , achieving a deep understanding of lengthy PDF documents remains a highly challenging task.
 To handle this task, we explore three approaches to input the PDF documents into these MLLMs: pure text content, page images, and parsed interleaved text-image content.
-For some baselines with plain text input modality, we also report results based on retrieval enhancement techniques  [45].
-The experimental results are reported in Tab.
-3 and Tab.
+For some baselines with plain text modality input, we also report the results based on Text RAG.
+For the pure vision solutions, we compared some document understanding models that can accept multi-page image input.
+To ensure a fair comparison regarding data availability, we not only compared against off-the-shelf RAG models but also fine-tuned the retrievers of the RAG baselines (BGE-M3  chen2024bge , VisRAG yu2024visrag ) on the PaperPDF dataset.
+Table 6: Comparison with commercial products (tested in Sep 2024) for long PDF understanding.
+The results are tested on a subset of 50 English PDFs.
+Note: These are the products rather than the models themselves.
+Model ANLS F1 Rouge GPT-Acc
+Gemini 1.5 pro gemini 26.6 29.0 29.8 67.9
+Kimi kimi4 28.5 33.6 31.1 74.7
+ChatGLM chatglm 31.2 35.4 32.0 73.5
+Qwen tongyi 36.0 40.3 35.5 78.1
+PDF-WuKong (ours) 41.8 43.2 40.7 77.5
+The experimental results for the open-source models are reported in Tab.
+5.
+The performance gap between PDF-WuKong and these fine-tuned baselines cannot be attributed solely to data exposure.
+Instead, it validates the architectural superiority of our end-to-end sparse sampling mechanism over disjoint retrieval-augmented pipelines.
+Generally, both document understanding models and RAG models oriented towards a single modality are inferior to our paradigm.
+Moreover, we compare PDF-WuKong with some closed-source commercial products in Tab.
+6, also showing our superiority.
+The qualitative comparisons between PDF-WuKong and these proprietary products on the long PDF understanding task are presented in Fig.
+5 (English) and Fig.
+6 (Chinese).
+Several key conclusions can be drawn from the results.
+1.
+Interleaved text-image content outperforms plain text input: When handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content.
+It is obvious that diagram information in the PDF plays a crucial role in document comprehension.
+This also indicates that PaperPDF places rigorous requirements on the visual information within documents.
+2.
+Interleaved text-image content outperforms pure vision input: When compared to pure visual input, interleaved text-image content provides superior performance.
+The main reason is the limited input resolution and number of tokens that the model can handle.
+3.
+Sampling enhances understanding: By introducing the single-modality RAG for MLLM to locate and sample evidence, their ability to understand long documents can be improved, which preliminarily verifies the feasibility of our motivation.
 4.
-Several key conclusions can be drawn from the results.
-Firstly, inputting parsed interleaved text-image content generally outperforms multiple page images.
-The main reason is the limited input resolution and number of tokens that the model can accept.
-Secondly, when handling tokens of similar scale, inputting parsed interleaved text-image content yields better performance compared to pure text content.
-It is obvious that diagram information in the PDF plays a crucial role for document comprehension.
-This also indicates that PaperPDF places rigorous requirements on the visual information within documents.
-Additionally, we observe that for the InternVL2 model, the approach of parsing interleaved text-image content as input outperforms InternVL2-RAG.
-However, the opposite conclusion is drawn for IXC2-VL.
-We hypothesize that this discrepancy may be due to the max length setting in IXC2-VL, which could cause the model to overlook some critical information.
-Finally, benefiting from the inclusion of the spare sampler, our proposed PDF-WuKong model not only surpasses the existing state-of-the-art open-source model InternVL2 by approximately 7% on both the Chinese and English subsets, but also demonstrates competitive performance comparable to proprietary products.
-Moreover, due to the integration of the sparse sampler, PDF-WuKong maintains efficiency in inference token cost.
-Model # param ANLS F1 ROUGE
-Qwen-VL [57] 9.6B 26.4 19.6 18.3
-Monkey [58] 9.8B 30.0 24.4 22.3
-mPLUG-Owl2 [25] 8.2B 19.5 20.3 22.7
-Emu2-Chat [59] 37B 26.0 24.4 23.4
-MiniCPM-2.5 [60] 8.5B 31.8 28.2 24.8
-IXC2-VL [49] 8B 23.4 20.8 21.3
-IXC2-4KHD [24] 8B 24.5 20.0 18.0
-CogVLM2 [61] 17B 24.8 27.4 26.3
-[…]
-[…]
+Limitations of single-modal sampling: Text-only RAG cannot focus on visual information, while visual-only RAG retrieves page images and cannot extract more fine-grained evidence.
+The full-page dense text images bring greater challenges to the image encoder.
+5.
+Superior performance on long PDF understanding: Benefiting from the inclusion of the multi-modal spare sampler, our proposed PDF-WuKong not only surpasses InternVL2 by approximately 7% on both the Chinese and English subsets but also demonstrates competitive performance compared with proprietary products.
+Moreover, due to the integration of the end-to-end sparse sampler, PDF-WuKong maintains efficiency during training and inference with fewer tokens.
+Figure 6: Examples of PDF-WuKong on Chinese documents. The red box indicates the evidence that the correct answer depends on.
 Table 7: Performance comparison with other DocVLMs for PDF multimodal understanding on the Single-Evidence Subset.
 $\dagger$ denotes the page image input, aligning with other models in the table; * indicates that our model utilizes parsed content as input.
-To compare with more open-source document MLLMs, considering that most of these models can only handle single-page documents, we construct a subset of the PaperPDF benchmark, only containing test samples with single evidence.
-Therefore, we provided all models with only one page containing the evidence as their input.
+Model # param ANLS F1 ROUGE
+Qwen-VL Qwen-VL 9.6B 26.4 19.6 18.3
+Monkey li2023monkey 9.8B 30.0 24.4 22.3
+mPLUG-Owl2 hu2024mplugdocowl2 8.2B 19.5 20.3 22.7
+Emu2-Chat sun2024generative 37B 26.0 24.4 23.4
+MiniCPM-2.5 hu2024minicpm 8.5B 31.8 28.2 24.8
+IXC2-VL internlmxcomposer2 8B 23.4 20.8 21.3
+IXC2-4KHD internlmxcomposer2_4khd 8B 24.5 20.0 18.0
+CogVLM2 hong2024cogvlm2 17B 24.8 27.4 26.3
+$\text{PDF-WuKong (ours)}^{\dagger}$ 8.5B 36.6 35.2 31.7
+$\text{PDF-WuKong (ours)}^{*}$ 8.5B 41.5 42.8 39.8
+The main reason for not using our data to train other MLLMs is that they cannot receive lengthy multimodal documents as input.
+Most models only support single-page input for training and few models can handle up to 8 or 20 pages.
+They are unable to be trained on our data.
+To compare with more open-source document MLLMs, we construct a subset of the PaperPDF benchmark, which only contains test samples with single evidence.
+This subset serves as a critical benchmark for fair comparison.
+It allows us to compare the fundamental multimodal understanding capabilities of our PDF-WuKong against other models under the same single-page input setting.
+Therefore, we feed all models with only one page containing the evidence as their input.
 The pages are input in the form of images.
@@
 As shown in Tab.
-5, our model’s capability on this subset is significantly better than other document models.
-Moreover, the document parsing-based paradigm used in our approach is superior to the purely visual paradigm.
+7, our model’s capability on this subset is significantly better than other document models.
+Moreover, introducing a parsed-input paradigm significantly reduces the difficulty of understanding text-rich documents, yielding approximately a 7% performance improvement on the Single-Evidence subset.
+This improvement can be primarily attributed to two factors.
+First, the parsed input allows the model to sample more fine-grained evidence for answer generation.
+Second, compared to raw full-page images, the parsed image data substantially relaxes the model’s dependency on high-resolution input, thereby enhancing both efficiency and robustness.
 5.3 Document-oriented VQA
-To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks and compare PDF-WuKong with other representative models.
-First, we evaluate the performance of PDF-WuKong on single-page document datasets [35, 37, 40].
+To validate the strong capability of our model in other document understanding scenarios, we conduct experiments on several public benchmarks.
+It is worth noting that while our primary pipeline utilizes parsed PDF content, PDF-WuKong is natively compatible with non-PDF inputs.
+In the following experiments, we feed the model with page images rather than parsed PDF blocks.
+PDF-WuKong treats each page as a visual block and can directly retrieve the pages containing the evidence.
+First, we evaluate the performance of PDF-WuKong on single-page document datasets Mathew_2021_WACV ; masry2022chartqa ; Mathew_2022_WACV .
 As shown in Tab.
-6, our model achieves comparable performance on single-page document understanding.
+8, our model achieves comparable performance on single-page document understanding.
 This demonstrates that PDF-WuKong can effectively handle various types of documents and questions, showcasing its versatility in document-oriented visual question-answering tasks.
-In addition, we assess the performance of traditional specialized models and MLLMs on two existing multi-page document QA datasets.
+Table 8: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks.
+Model DocVQA ChartQA InfoVQA
+Qwen-VL Qwen-VL 65.1 65.7 35.4
+Monkey li2023monkey 66.5 65.1 36.1
+Text-Monkey liu2024textmonkey 73.0 66.9 28.6
+MiniCPM-V-2.5 hu2024minicpm 84.8 - -
+Vary-base wei2025vary 76.3 66.1 -
+TextHawk yu2024texthawk 76.4 66.6 50.6
+IXC2-4KHD-16 internlmxcomposer2_4khd 84.9 80.1 60.8
+DocOwl 2 hu2024mplugdocowl2 80.7 70.0 46.4
+CREAM zhang2024cream 79.4 - 53.6
+PDF-WuKong (ours) 85.1 80.0 61.3
+Table 9: Performance comparison with other DocVLMs for multi-page document understanding.
+Model MP-DocVQA DUDE
+LayoutLMv3 huang2022layoutlmv3 55.1 20.3
+Longformer beltagy2020longformer 55.1 27.1
+BigBird zaheer2020big 58.5 26.3
+Hi-VT5 TITO2023109834 61.8 35.7
+DocFormerv2 appalaraju2024docformerv2 76.4 48.4
+GRAM blau2024gram 83.0 53.4
+GPT-4V (2024-06) 2023gpt4V - 53.9
+Idefics3-8B laurenccon2024obelics 67.2 38.7
+DocOwl2 hu2024mplugdocowl2 69.4 46.7
+CREAM zhang2024cream 65.3 52.5
+SV-RAG chen2025svrag 71.0 45.0
+PDF-WuKong (ours) 76.9 56.1
+In addition, we assess the performance of specialized models and MLLMs on two existing multi-page document QA datasets.
 The results shown in Tab.
-7 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 [25].
-Notably, on complex multi-page document datasets like DUDE [17], PDF-WuKong outperforms GPT-4V [44].
-This improvement is attributed to our sparse sampler, which effectively filters out useful information from multi-page documents, enabling the model to focus on relevant content.
-Furthermore, we conduct zero-shot evaluations on a new long multimodal document understanding benchmark MM-NIAH [18].
+9 indicate that our model’s performance in multi-page document scenarios is comparable to these specialized models and far surpasses the latest document MLLM DocOwl2 hu2024mplugdocowl2 .
+Notably, on complex multi-page document datasets like DUDE van2023document , PDF-WuKong outperforms GPT-4V 2023gpt4V .
+This improvement is attributed to the sparse sampler, which extracts useful information from multi-page documents, enabling the model to focus on relevant content.
+Furthermore, we conduct zero-shot evaluations on a long multimodal document understanding benchmark MM-NIAH wang2024needle .
 As shown in Tab.
-8, our model uses the fewest parameters yet achieves the second-best performance.
-Although InternVL-V1-5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5 billion more parameters than our model.
+10, our model uses the fewest parameters yet achieves the second-best performance.
+Although InternVL-V1.5-RAG surpasses PDF-WuKong by 2.8%, it utilizes 36.5B more parameters than our model.
 Moreover, as the context length of the multimodal documents increases, the performance of our model remains stable, while that of other models significantly decreases.
 At a context length of 64K, PDF-WuKong even achieves the best performance, demonstrating its robustness in handling long-context multimodal inputs.
-Model DocVQA ChartQA InfoVQA
-Qwen-VL [57] 65.1 65.7 35.4
-Monkey [58] 66.5 65.1 36.1
-Text-Monkey [10] 73.0 66.9 28.6
-MiniCPM-V-2.5 [60] 84.8 - -
-Vary-base [12] 76.3 66.1 -
-TextHawk [62] 76.4 66.6 50.6
-IXC2-4KHD-16 [24] 84.9 80.1 60.8
-DocOwl 2 [25] 80.7 70.0 46.4
-CREAM [26] 79.4 - 53.6
-PDF-WuKong (ours) 85.1 80.0 61.3
-Table 6: Performance comparison with other DocVLMs on single-page document-oriented VQA benchmarks.
-Model MP-DocVQA DUDE
-LayoutLMv3 [54] 55.1 20.3
-Longformer [55] 55.1 27.1
-BigBird [56] 58.5 26.3
-Hi-VT5 [46] 61.8 35.7
-DocFormerv2 [63] 76.4 48.4
-GRAM [16] 83.0 53.4
-GPT-4V (2024-06) [44] - 53.9
-Idefics3-8B [64] 67.2 38.7
-DocOwl2 [25] 69.4 46.7
-CREAM [26] 65.3 52.5
-PDF-WuKong (ours) 76.9 56.1
-Table 7: Performance comparison with other DocVLMs for multi-page document understanding.
+Table 10: Performance comparison with other DocVLMs on $M$-NIAH. The evaluation approach aligns with the benchmark.
 Model #param Overall 1K 4K 16K 64K
-Emu2-Chat [59] 37B 8.8 38.9 18.2 0.0 0.0
-VILA1.0-13b [65] 13B 15.7 41.9 33.2 8.6 0.1
-llava-v1.6-13b [23] 13B 16.9 43.7 34.9 13.6 0.0
-llava-v1.6-34b [23] 34B 20.6 57.4 45.1 8.2 0.0
-InternVL1.5 [11] 26B 41.1 59.5 50.1 41.9 16.6
-InternVL1.5-RAG [11] 45B 46.1 59.5 50.1 44.9 39.3
+Emu2-Chat sun2024generative 37B 8.8 38.9 18.2 0.0 0.0
+VILA1.0-13b lin2024vila 13B 15.7 41.9 33.2 8.6 0.1
+llava-v1.6-13b liu2024llavanext 13B 16.9 43.7 34.9 13.6 0.0
+llava-v1.6-34b liu2024llavanext 34B 20.6 57.4 45.1 8.2 0.0
+InternVL1.5 chen2024far 26B 41.1 59.5 50.1 41.9 16.6
+InternVL1.5-RAG chen2024far 45B 46.1 59.5 50.1 44.9 39.3
 PDF-WuKong (ours) 8.5B 43.3 53.0 43.9 43.0 42.1
-Table 8: Performance comparison with other DocVLMs on MM-NIAH. The evaluation approach aligns with the benchmark.
+5.4 Fine-grained Evaluation on Long PDF Understanding
+To comprehensively assess the capabilities of PDF-WuKong in long-document processing, we conducted a fine-grained evaluation on the PaperPDF benchmark, as presented in Table 11.
+We can observe that the model exhibits superior performance in figure and table understanding tasks compared to other categories questions, validating that the sparse sampling mechanism effectively captures critical visual evidence within long documents.
+However, formula understanding remains weaker, which may be attributed to the loss of 2D structural semantics during text linearization and the scarcity of relevant training data.
+Additionally, the model performs slightly better on Multi-hop tasks, a trend suggesting that the model potentially benefits from utilizing global redundant context, unlike single-hop tasks that focus on locating isolated details.
+Table 11: Performance of PDF-WuKong across different question types.
+TU, FU, FoU, MuU, MuL, SiU and SiL represent Table Understanding, Figure Understanding, Formula Understanding, Multi-hop Understanding, Multi-hop Logical Reasoning, Single-hop Understanding and Single-hop Logical Reasoning, respectively.
+Question PDF-WuKong InternVL2
+Type ANLS F1 ROUGE ANLS F1 ROUGE
+TU 44.15 46.32 45.09 40.65 45.44 45.71
+FU 46.05 48.38 46.46 40.82 44.68 45.03
+FoU 39.67 39.81 38.08 26.67 28.59 30.72
+MuU 44.07 45.30 45.01 35.83 37.07 39.13
+MuL 41.13 42.40 38.92 32.13 33.37 34.26
+SiU 42.32 43.33 40.30 28.69 31.11 30.90
+SiL 39.05 38.41 35.55 27.04 28.16 28.46
 5.5 Ablation Study
 To comprehensively evaluate the effectiveness of our contributions, we conduct ablation studies focusing on the impact of the sparse sampler, the dataset, the document length, and sampling strategies.
-These experiments are based on English PaperPDF for training and evaluation to avoid interference from other factors.
+These experiments are based on the English subset of our PaperPDF for training and evaluation to avoid interference from other factors.
 Sparse sampler
@@
 Introducing the sparse sampler significantly improved the model’s accuracy, as evidenced in Tab.
-9, by efficiently selecting the most relevant content for each query.
+12, by efficiently selecting the most relevant content for each query.
 Furthermore, end-to-end joint training of the sparse sampler and the MLLM led to additional performance gains compared to training them separately.
 This indicates that our end-to-end optimization of multimodal representation and question answering can further promote the document understanding ability of MLLM.
+Table 12: Ablation study on the impact of sparse sampler
 Sparse Sampler End-to-End ANLS F1 ROUGE
@@
 42.6 43.6 40.2
-Table 9: Ablation study on the impact of sparse sampler
 Dataset
-We retrain PDF-WuKong on various subsets of our English PaperPDF, and the results are shown in Tab.
-10.
-Increasing the amount of training data leads to consistent improvements in the model’s accuracy, proving that our dataset follows scaling laws.
-In addition, we verify the effectiveness of our two data construction methods.
+We retrain PDF-WuKong on various dataset settings of the English subset of our PaperPDF, and the results are shown in Tab.
+13.
+Increasing the amount of training data leads to consistent improvements, proving that our dataset is of high quality and follows scaling laws.
+Besides, we verify the effectiveness of our two data construction methods.
 Under the same data scale, multi-evidence data can enhance the model’s complex reasoning ability.
+Table 13: Ablation study on dataset setting
 Dataset ANLS F1 ROUGE
@@
 Multi (100k) + Single (100k) 40.0 41.6 38.9
-Table 10: Ablation study on dataset setting
 Document length
@@
 Results in Fig.
-4 demonstrate that our model’s performance and token counts remain relatively stable across documents of varying lengths.
-This stability indicates that the sparse sampler effectively reduces the input size to a reasonable level, regardless of the original document length.
-In contrast, the baseline MLLM without the sparse sampler is unable to handle long documents effectively.
+7 show that our model’s performance and token counts remain stable across documents of varying lengths.
+This stability indicates that the sparse sampler effectively reduces the input size, regardless of the original document length.
+In contrast, the baseline without the sparse sampler is unable to handle long documents effectively.
 Its performance deteriorates significantly as the document length increases.
@@
 Figure 7: Ablation study of different document length
+Table 14: Ablation study of different sampling strategy. P and R, represent the Precision and Recall of the evidence sampling module.
+Sampling Retrieval QA Token
+chunks P R ANLS F1 ROUGE
+Top 1 46.00 45.60 39.20 38.28 35.26 1186
+Top 3 22.53 66.72 42.09 43.06 39.69 1452
+Top 5 14.96 73.46 42.59 43.63 40.19 1789
+Top 10 8.40 81.79 43.01 44.22 40.67 2386
+Top 15 5.94 86.61 43.19 44.57 42.08 2704
+Top 20 4.63 89.64 43.42 45.02 42.30 3364
 Sampling strategy
-We explore the impact of different numbers of text blocks or diagrams selected by the sparse sampler.
+To investigate the effectiveness of the sparse sampler in evidence location, we conduct an ablation study on the number of selected blocks.
+Beyond standard QA performance evaluation, we further explore the precision and recall of the sampled evidence , following the previous studies souibgui2025docvxqa ; chen2025towards .
 As shown in Tab.
-11, setting a small top $k$ can lead to missing crucial information needed for accurately answering queries, thus reducing performance.
-Conversely, a larger top $k$ introduces redundant information and increases computational costs without significantly enhancing accuracy.
-To strike a balance between performance and resource efficiency, we use the top 5 as the default setting.
-Sampling chunks ANLS F1 ROUGE Tokens
-Top 1 39.20 38.28 35.26 1186
-Top 3 42.09 43.06 39.69 1452
-Top 5 42.59 43.63 40.19 1789
-Top 10 43.01 44.22 40.67 2386
-Top 15 43.19 44.57 42.08 2704
-Top 20 43.42 45.02 42.30 3364
-Table 11: Ablation study of different sampling strategy
+14, setting a small top $k$ results in relatively low recall, which implies that the crucial evidence required for accurately answering queries is often omitted, thereby constraining the model’s performance.
+Conversely, although a larger top-$k$ indeed improves evidence recall, it comes at the cost of a sharp decline in precision, introducing substantial non-evidential redundancy that interferes with the model’s reasoning.
+As a result, when top-$k$ exceeds 5, the model exhibits diminishing returns, where a substantial increase in token consumption yields only marginal gains in QA accuracy.
+To strike an optimal balance between performance and resource efficiency, we adopt Top-5 as the default setting.
 6 Conclusion
@@
 Our approach maintains high accuracy and efficiency even as document length increases, offering a scalable and interpretable solution for practical applications in document understanding.
-Appendix A Algorithm
-Algorithm 1 shows the detailed inference process of PDF-WuKong.
-The training pipeline is shown in Algorithm 2.
-Our PDF-WuKong can achieve efficient and accurate understanding of long PDFs with end-to-end sparse sampling.
-Algorithm 1 Inference pipeline for PDF-WuKong
-1:Input: PDF document $D$, user query $q$
-2:Output: Generated answer $a$
-3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
-4:Stage 1: Document Parsing
-5:Parse the input document $D$ into text blocks and images:
-$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
-6:Stage 2: Sparse Sampling
-7:Encode all text blocks and images and cache all candidate vector embeddings:
-$E_{T}=\{e_{T_{1}},e_{T_{2}},\dots,e_{T_{n}}\}\leftarrow\text{En\_T}(\{T_{1},T_{2},\dots,T_{n}\}),$
-$E_{I}=\{e_{I_{1}},e_{I_{2}},\dots,e_{I_{m}}\}\leftarrow\text{En\_I}(\{I_{1},I_{2},\dots,I_{m}\})$
-8:Encode the user query $q$:
-$e_{q}\leftarrow\text{En\_T}(q)$
-9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{E_{T},E_{I}\}$:
-$S_{T}=\{\text{Sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
-$S_{I}=\{\text{Sim}(e_{q},e_{I_{j}})\mid j=1,2,\dots,m\}$
-10:Select the top-$k$ relevant text blocks and images:
-$(T,I)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$
-11:Stage 3: Answer Generation
-12:Input the query $q$ and the selected tokens into the LLM:
-$a\leftarrow\text{LLM}(q,(T,E_{I})_{top})$
-13:Return the generated answer $a$.
-Appendix B Visualization Results
-B.1 Qualitative Comparison with Other Products
-The qualitative comparison between PDF-WuKong and other proprietary products on the long PDF understanding task is presented in Fig.
-5.
-PDF-WuKong achieves significant performance advantages over others such as Qwen [66], ChatGLM [67], Kimi [68], and Gemini-pro [69].
-The success stems from its ability to effectively sample key evidence relevant to the question and its strong multimodal understanding capabilities.
-This observation further validates the effectiveness of our method in enhancing the long PDF understanding capabilities of MLLMs.
-Algorithm 2 Training pipeline for PDF-WuKong
-1:Input: PDF document $D$, user query $q$, ground truth answer $gt$
-2:Output: Final loss function $\mathcal{L}_{\text{total}}$
-3:Initialize: Text encoder En_T, image encoder En_I, large language model LLM
-4:Stage 1: Data Preparing
-5:Text blocks and images:
-$\{T_{1},T_{2},\dots,T_{n}\},\{I_{1},I_{2},\dots,I_{m}\}\leftarrow\text{Parser}(D)$
-6:Stage 2: Multimodal encoding
-7:Encode the user query, positive and negative samples:
-$e_{q}\leftarrow\text{En\_T}(q)$
-$E_{T}=\{e_{T_{P}},e_{T_{N}}\}\leftarrow\text{En\_T}(\{T_{P},T_{N}\}),$
-$E_{I}=\{e_{I_{P}},e_{I_{N}}\}\leftarrow\text{En\_I}(\{I_{P},I_{N}\}),$
-8:Calculate the contrastive learning loss:
-$\mathcal{L}_{\text{rep}}(e_{q},\{e_{T_{P}},e_{I_{P}}\},\{e_{T_{N}},e_{I_{N}}\})$
-9:Stage 3: Output prediction of MLLM
-10:Input the query, the positive text, and the positive image tokens from the shared image encoder En_I:
-$a\leftarrow\text{LLM}(q,T_{P},e_{I_{P}})$
-11:Calculate the cross-entropy loss:
-$\mathcal{L}_{\text{QA}}(a,gt)$
-12:Stage 4: Optimize model in an end-to-end manner
-13:Update model parameters according to the joint loss:
-$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rep}}+\mathcal{L}_{\text{QA}}$
-14:Return the final loss $\mathcal{L}_{\text{total}}$.
-Figure 5: Qualitative comparison between PDF-WuKong with other proprietary products. The red box indicates the evidence that the correct answer depends on.
-B.2 Visualization Results on Chinese Documents
-In addition to the outstanding performance on English documents, PDF-WuKong also demonstrates remarkable capabilities on Chinese documents.
+7 Limitation
+It is important to note that unlike the manually verified test set, the training set is constructed via an automated pipeline.
+Consequently, it may contain a certain degree of noise stemming from PDF parsing errors (e.g., imperfect layout analysis in complex documents) or alignment inconsistencies in multi-evidence extraction.
+While this noise is typical in large-scale pre-training datasets, users should be aware of these potential imperfections.
+For our proposed dataset, current documents are mainly limited to academic papers, so the layout format and subject matter of the documents are relatively simple.
+We will expand the dataset with more diverse documents.
+Besides, our model is not specifically designed for some global queries, which will become a key research problem in the future.
+Data Availability Statement
+The datasets used during and/or analysed in the current study are available in our PaperPDF repository [Link], the DocVQA repository [Link], the ChartQA repository [Link], the InfoVQA repository [Link], the MP-DocVQA repository [Link], the DUDE repository [Link], and the $M$-NIAH repository [Link].
+Declaration
+The authors have no relevant financial or non-financial interests to disclose.
+Acknowledgement
+This research is supported by the National Science Fund for Distinguished Young Scholars of China (Grant No.62225603) and the National Natural Science Foundation of China (Grant No. 62576147).
+Appendix A Prompts and Examples for PaperPDF Dataset
 Fig.
-6 shows its qualitative performance on the Chinese long PDF document understanding task.
-It effectively samples relevant evidence within the PDF and demonstrates the ability to derive accurate answers based on the sampled evidence.
-Appendix C More Dataset Details
-The Q-E-A triplets in the PaperPDF dataset are derived from approximately 60k English and 10k Chinese documents, encompassing nearly 70 disciplines such as computer science, engineering, and materials science.
-To comprehensively illustrate the characteristics of the PaperPDF dataset, we conduct a detailed statistical analysis.
-The key statistics of the source documents in PaperPDF are presented in Tab.
-12.
-Compared to previous multi-page document datasets, PaperPDF not only includes a significantly larger number of documents but also features more pages and OCR tokens.
-More importantly, PaperPDF encompasses both Chinese and English documents with interleaved text and images.
-Our dataset provides an important innovation driver and a comprehensive benchmark for the development of the multimodal long PDF document understanding task.
-For the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence, Fig.
 8 - Fig.
-11 show their prompt engineering and a corresponding example, respectively.
-Figure 6: Examples of PDF-WuKong on Chinese documents. The red box indicates the evidence that the correct answer depends on.
-MPdataset Document Language Page Image Token
-DUDE [17] 5k En 6 - 1831
-MP-Docvqa [46] 6k En 8 - 2026
-PaperPDF (En) 60k En 25 12 11371
-PaperPDF (Zh) 10k Zh 11 7 3413
-PaperPDF 70k En+Zh 23 11 10234
-Table 12: Statistics of the source documents, where Page, Image, and Token represent the average number of pages per document, the average number of images in the document, and the average number of OCR tokens per document, respectively.
+12 show the prompt engineering for constructing the five types of data (Text-only, Image-only, Image-text, Section, Cross-paragraph) in single evidence and multiple evidence.
+There are also corresponding examples for these five types.
 Text-only Question Generation Prompt Task Definition: Create 2 academic questions from a given research paper paragraph.
@@
 Figure 12: Cross-paragraph Q-E-A triplets generation prompt and data example.
-Text-only Data Example Query: What is the impact of using the same dataset for optimizing and measuring the performance of a model?
-Text: Here, again, the unfair advantage of optimizing (selecting the models for the ensemble) and measuring performance on the same dataset appears.The advantage is small but systematic for the test split of ISIC (Fig.
-5a); it is much more apparent for the challenging collection of clinical images of EDRA Atlas (Fig.
-5b).
-Answer 1: It can lead to an unfair advantage for the model.
-Answer 2: Optimizing a model involves selecting certain parameters or features that improve its performance on a given dataset.
-If the same dataset is used to measure the model’s performance, it may lead to an unfair advantage as the model has already been “tuned” to that specific dataset.
-Text-image Data Example Query: What is the best performing method for both detection and classification, according to the provided figure?
-Text: 3) The last one is our proposed SFCN-OPI with both sibling branches and OPI (Ours in Table 1).
-Figure: Table 1[Uncaptioned image] Caption: Experimental results of ablation analysis, different training strategies of our method and comparison with other approaches.
-Note: the -means the results were not reported by that method.
-Answer 1: Ours.
-Answer 2: The best performing method for both detection and classification is Ours, as it achieves the highest F1 scores for both tasks.
-This can be seen in the “Ours” row of the table, where the F1 score for detection is 0.834 and the F1 score for classification is 0.742.
-Appendix D Limitation
-For our proposed dataset, current documents are mainly limited to academic papers, so the layout format and subject matter of the documents are relatively simple.
-We will expand the dataset with more diverse documents.
-Besides, our model is not specifically designed for some global queries, which will become a key research problem in the future.